### 19 Mar 2026 v1

# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

[…]
On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization
[…]
On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization
[…]
Motivated by the principle of satisficing in decision-making, we study satisficing regret guarantees for nonstationary $K$-armed bandits. We show that in the general realizable, piecewise-stationary setting with $L$ stationary segments, the optimal regret is $\Theta(L\log T)$ as long as $L\geq 2$. This stands in sharp contrast to the case of $L=1$ (i.e., the stationary setting), where a $T$-independent $\Theta(1)$ satisficing regret is achievable under realizability. In other words, the optimal regret has to scale with $T$ even if just a little nonstationarity presents. A key ingredient in our analysis is a novel Fano-based framework tailored to nonstationary bandits via a post-interaction reference construction. This framework strictly extends the classical Fano method for passive estimation as well as recent interactive Fano techniques for stationary bandits. As a complement, we also discuss a special regime in which constant satisficing regret is again possible.
[…]
nonstationary bandits, satisficing regret, fano methods, post-interaction reference
1 Introduction
The multi-armed bandit (MAB) problem is a classical framework for studying the exploration–exploitation trade-off in sequential decision-making. In the standard MAB setting, a learner repeatedly chooses among $K$ arms with unknown, arm-specific stationary reward distributions, with the goal of minimizing cumulative regret (Robbins 1952, Lai and Robbins 1985). For the canonical MAB problem and its many variants, a rich literature has established sharp performance guarantees; see, for example, Lattimore and Szepesvári (2020).
While identifying the optimal action is essential in some high-stakes settings, many practical applications only require a “good-enough,” or satisficing, decision. Introduced by Simon (1956), satisficing aims for an acceptable solution rather than the absolute best one. This perspective is common in practice. For instance, recommendation and advertising systems often deploy a candidate only if its KPI exceeds a threshold $S$ with sufficient confidence (Kohavi et al. 2020, Xu et al. 2015), and inventory control typically targets a desired service level (e.g., 90–95%) to mitigate stockout costs (Fitzsimons 2000, Jing and Lewis 2011).
These considerations motivate the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2022, Rouyer et al. 2024, Feng et al. 2024, Hajiabolhassan and Ortner 2025), where the learner aims to pull a satisficing arm—an arm whose mean reward is no less than a prescribed threshold SS—as often as possible. Performance is measured by the satisficing regret, defined as the cumulative shortfall of the selected arm’s mean reward relative to $S$. To the best of our knowledge, existing results on satisficing regret primarily focus on stationary environments: Tamatsukuri and Takahashi (2019), Michel et al. (2022), Rouyer et al. (2024) study the $K$-armed bandit setting, Feng et al. (2024) extend the framework to more general bandit optimization problems, and Hajiabolhassan and Ortner (2025) study the average-reward Markov decision process setting. Under a realizability condition—namely, that at least one (initially) unknown arm is strictly satisficing with mean exceeding SS—these works show that the satisficing regret can be bounded by a constant independent of the decision horizon $T$.
In many real-world scenarios, however, stationarity fails: consumer preferences drift with seasons and trends (Ardeshiri et al. 2019), and platform marketplaces (e.g., ride-hailing) experience demand surges and changing competitive conditions (Battifarano and Qian 2019). As a result, an arm’s reward distribution may evolve over time. Despite the growing literature on satisficing regret in stationary environments, much less is understood in the nonstationary setting that frequently arises in practice. Motivated by this gap, we study satisficing regret minimization under nonstationarity, focusing on piecewise-stationary bandits with $L$ stationary segments. Building on the stationary case (corresponding to $L=1$), it is natural to ask:
Can we similarly achieve a $T$-independent satisficing regret in the piecewise-stationary setting?
At first glance, this appears plausible since if the change points were known in advance, one could simply restart a stationary satisficing algorithm on each segment, leading to an $\mathcal{O}(L)$ satisficing regret.
1.1 Our Contributions
In this work, however, we provide a negative answer to the above conjecture. Specifically,
  • In Section 4, we establish an $\Omega\bigl(L\log(T/L)\bigr)$ lower bound on the satisficing regret for nonstationary $K$-armed bandits. We then propose Algorithm 1, which achieves an almost matching upper bound of $\mathcal{O}(L\log T)$. Together, these results show that the additional $\log T$ factor is unavoidable in general when the change points are unknown, thereby ruling out the conjectured $\mathcal{O}(L)$ satisficing regret guarantee.
  • To establish the lower bounds above, we extend recent interactive Fano techniques for bandit regret lower bounds (Chen et al. 2024, Shufaro et al. 2024) from stationary to nonstationary environments. More importantly, we introduce a post-interaction reference framework that strictly generalizes the classical Fano approach and is particularly well suited to nonstationary bandits. The key feature is that an auxiliary random variable (the “reference”) may influence the data-generating process and is revealed to the estimator after the interaction. This leads to two complementary principles for lower-bounding regret (Section 3.2): (i) reducing a portion of the regret to an identification error under a suitable separation condition, and (ii) interpreting a portion of the regret as an information cost. We further emphasize that this framework may extend to broader classes of nonstationary bandit problems and may be of independent interest.
  • In Section 5, we complement the general negative result by identifying a special case in which constant satisficing regret is still achievable. Specifically, when there exists an always-realizable arm and no satisficing arm ever becomes non-satisficing, the stationary algorithms of Michel et al. (2022) can be applied directly in the nonstationary setting, yielding constant satisficing regret that is independent of both the horizon $T$ and the number of stationary segments $L$.
1.2 Related Work
Nonstationary bandits.
Most work on nonstationary bandits evaluates performance through dynamic regret, which compares the learner with the best arm at each round. For piecewise-stationary environments with at most $L$ segments, sliding-window and change-point-based methods achieve $\widetilde{\mathcal{O}}(\sqrt{LT})$ dynamic regret (Garivier and Moulines 2011, Auer et al. 2019). Under a variation-budget model with total drift $D$, the optimal rate is $\widetilde{\mathcal{O}}(D^{1/3}T^{2/3})$ (Besbes et al. 2014, Cheung et al. 2019), and Wei and Luo (2021) obtain the unified guarantee $\widetilde{\mathcal{O}}(\min\{\sqrt{LT},D^{1/3}T^{2/3}\})$. These results address a stronger benchmark than ours: dynamic regret competes with the instantaneous best arm, whereas our objective only requires selecting an arm whose mean exceeds a fixed threshold.
Stationary satisficing regret.
Our work is most closely related to the literature on threshold-based satisficing regret in stationary bandits. Michel et al. (2022), Rouyer et al. (2024), and Feng et al. (2024) study cumulative satisficing regret under a fixed threshold and show that, under realizability, constant regret is achievable in stationary settings. Our paper shows that this phenomenon changes sharply under nonstationarity: once the environment may change at unknown times, constant satisficing regret is no longer achievable in general.
Threshold-based pure exploration and other satisficing notions.
In stationary settings, threshold-based objectives are also studied in pure exploration under the name thresholding bandits (Locatelli et al. 2016, Mukherjee et al. 2017). These works aim to identify the arms whose means exceed a threshold and are evaluated by probability of error or sample complexity, rather than cumulative satisficing regret. Although such methods can identify a satisficing arm with high probability, a naive explore–then–exploit reduction is not sufficient to obtain optimal cumulative satisficing regret even in stationary problems (Michel et al. 2022, Feng et al. 2024).
Russo and Van Roy (2018) introduce a discounted notion of satisficing in stationary bandits, motivated by time-sensitive learning. Their setting and benchmark differ from ours: they study discounted regret rather than cumulative threshold shortfall.
For nonstationary bandits, as highlighted in Liu et al. (2023b, a), the standard notion of dynamic regret can steer algorithm design toward overly aggressive exploration; it can therefore be more meaningful to compete against a satisficing benchmark. Min and Russo (2023) study nonstationary bandits in a Bayesian model and introduce a notion of satisficing based on competing with a satisficing action sequence $A^{\dagger}$ (Min and Russo 2023, Sec. 7). Their ‘satisficing regret” is defined as regret relative to $A^{\dagger}$. In contrast, we study a threshold-based notion of satisficing regret, defined as the cumulative shortfall from a fixed threshold $S$. Our formulation does not require that any arm achieves mean exactly $S$ at each time. Even if one sets $A^{\dagger}$ to be a realizable arm sequence, the two regret notions are not directly comparable: regret relative to $A^{\dagger}$ can be negative at times when the learner selects an arm with strictly larger mean than $A_{t}^{\dagger}$, whereas our satisficing regret is nonnegative by definition. Finally, Min and Russo (2023) establishes asymptotic lower bounds for the classical dynamic regret, while we provide minimax-optimal finite-time guarantees for threshold-based satisficing regret.
1.3 Notation
For integers $M\geq 1$, define $M]:=\{1,2,\ldots,M\}$. For integers $0\leq N\leq M$, define $N,M]:=\{N,N+1,\ldots,M\}$. For a buffer $\mathcal{B}=(b_{1},\ldots,b_{m})$ of length $m$, we write $\mathcal{B}[s]:=b_{s}$ for its $s$-th entry, $s\in[m]$. For a set $\mathcal{A}$, we write $|\mathcal{A}|$ for its cardinality, and $\mathrm{Unif}(\mathcal{A})$ for the uniform distribution on $\mathcal{A}$. For random variables $(X,Y)$ with joint law $P_{X,Y}$, the mutual information is
$I(X;Y):=\mathbb{E}_{X}\left[D_{\mathrm{KL}}\bigl(P_{Y\mid X}\,\|\,P_{Y}\bigr)\right],$ (1)
[…]
2 Problem Formulation
Nonstationary Bandits
We study a nonstationary $K$-armed bandit over a horizon $T\in\mathbb{N}$, with $K\lesssim T$. At each round $t\in[T]$, each arm $a\in[K]$ has an unknown mean reward $\mu_{t}(a)\in[0,1]$. For integers $1\leq s\leq t\leq T$, define the number of stationary segments on $[s,t]$ as
$L[s,t]:=1+\sum_{j=s}^{t-1}\mathds{1}\left\{\exists a\in[K]\text{ such that }\ \mu_{j}(a)\neq\mu_{j+1}(a)\right\}.$
Let $L:=L[1,T]$ denote the total number of stationary segments, and let
$1=T_{0}<T_{1}<\cdots<T_{L}=T+1$
be the associated change-points that partition the horizon into $L$ stationary segments, where segment $\ell\in\{0,\dots,L-1\}$ is the interval $T_{\ell},\,T_{\ell+1}-1]$. A (possibly randomized) policy $\pi$ selects an action $a_{t}\in[K]$ at time $t$, adapted to the natural filtration $\mathcal{F}_{t-1}:=\sigma(a_{1},r_{1},\dots,a_{t-1},r_{t-1})$. Let $\Pi$ denote the class of all such policies. Conditional on $a_{t}$, the learner observes a reward $r_{t}=\mu_{t}(a_{t})+\xi_{t},$ where $(\xi_{t})_{t\geq 1}$ are independent, zero-mean, $1$-sub-Gaussian noises.
Satisficing Regret.
The decision maker specifies a threshold $S\in\mathbb{R}$, and performance is judged by whether the chosen arm meets this level. An arm $a$ is satisficing at time $t$ if $\mu_{t}(a)\geq S$. The one-step satisficing regret is $(S-\mu_{t}(a_{t}))_{+}:=\max\{0,S-\mu_{t}(a_{t})\}$, and the cumulative satisficing regret under policy $\pi$ is
$R_{T}^{S}(\pi;\nu):=\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\Big],$ (2)
where $\nu$ denotes the underlying nonstationary environment and $\mathbb{E}_{\pi,\nu}$ is over the policy’s randomness and the observation noise.
Satisficing nonstationary bandits.
For each segment $\ell\in\{0,\dots,L-1\}$, let $\mu_{a,\ell}:=\mu_{T_{\ell}}(a)$ denote the mean reward of arm $a\in[K]$ on segment $\ell$, and define the segment-wise optimal mean $\mu_{\ell}^{*}:=\max_{a\in[K]}\mu_{a,\ell}$. Throughout, we impose the following realizability condition.
Assumption 1 (Realizability).
For every segment $\ell\in\{0,\dots,L-1\}$, there exists at least one arm with mean strictly above the threshold, i.e., $\mu_{\ell}^{*}>S$.
Assumption 1 is standard in studies of satisficing regret in the stationary setting (Michel et al. 2022, Feng et al. 2024). We refer to the nonstationary bandit problem under Assumption 1, together with the objective (2), as the nonstationary satisficing bandit problem.
3 A Fano-Based Framework for Nonstationary Bandit Problems
Fano’s lemma (Reza 1994, Theorem 2.10) is a classical tool for minimax lower bounds in multi-hypothesis testing. Let $\mathcal{V}$ be a finite index set, let $\nu\sim\mathrm{Unif}(\mathcal{V})$ be a latent instance index, and let $Z$ denote the data generated by an algorithm interacting with the instance indexed by $\nu$. Fano’s lemma lower bounds the best achievable identification error in terms of the mutual information $I(\nu;Z)$:
$\inf_{\widehat{\nu}(Z)}\mathbb{P}\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\geq 1-\frac{I(\nu;Z)+\log 2}{\log|\mathcal{V}|}.$ (3)
Recently, building on (3), Chen et al. (2024), Shufaro et al. (2024) developed general Fano-type frameworks for proving regret lower bounds in stationary bandit problems. In Section 3.1, we review how these works leverage (3) to derive interactive Fano methods for stationary bandits.
Conditional Fano.
It is often useful to apply Fano’s lemma in a structured setting where the index takes the form $\nu=(u,b)\in\mathcal{U}\times\mathcal{B}$. Here the “reference” variable $b$ is independent of $u$, the observation $Z$ may depend on $b$, and $b$ is revealed to the estimator before it outputs $\widehat{u}$. A conditional form of Fano’s inequality (Scarlett and Cevher 2019, Theorem 3) then follows directly from (3):
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{I(u;Z\mid b)+\log 2}{\log|\mathcal{U}|}.$ (4)
In Section 3.2, building on (4), we develop a post-interaction reference framework which is particularly convenient for nonstationary bandits and yields two complementary principles for deriving regret lower bounds.
3.1 Classical Fano Reduction: Small Regret Implies Correct Identification
Draw an environment index $\nu\sim\mathrm{Unif}(\mathcal{V})$. Fix any policy $\pi$, and let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ denote the full interaction transcript obtained by running $\pi$ for $T$ rounds in environment $\nu$. Let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$, and let $\bar{P}:=\frac{1}{|\mathcal{V}|}\sum_{\tau\in\mathcal{V}}P_{\tau}$ be the resulting mixture distribution. By the definition of mutual information and the convexity of $D_{\mathrm{KL}}(P\|\cdot)$ in its second argument (Cover 1999), we have
$I(\nu;Z)=\frac{1}{|\mathcal{V}|}\sum_{\tau\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|\bar{P}\right)\leq\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right).$ (5)
Combining (5) with (3) yields
$\inf_{\widehat{\nu}(Z)}\mathbb{P}\!\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\geq 1-\frac{\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right)+\log 2}{\log|\mathcal{V}|}.$ (6)
In stationary bandits, a standard route to regret lower bounds is to reduce small regret to accurate identification of the underlying instance. Concretely, one constructs an instance family $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ satisfying a separation condition at level $\Delta$ (see, e.g., Shufaro et al. 2024, Prop. 3.2 or the more general formulation in Chen et al. 2024, Eq. (10)), which enables the construction of an estimator $\widehat{\nu}(Z)$ from the transcript $Z$ such that
$\mathbb{P}\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\leq\mathbb{P}\bigl(\mathrm{Reg}_{T}(\pi,\nu)\geq T\Delta\bigl),$ (7)
where $\mathrm{Reg}_{T}$ denotes the regret of interest. By Markov’s inequality and (6), the Bayesian regret under the uniform prior satisfies
$\mathbb{E}\!\left[\mathrm{Reg}_{T}(\pi,\nu)\right]\;\geq\;T\Delta\cdot\Bigl(1-\frac{\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\!\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right)+\log 2}{\log|\mathcal{V}|}\Bigr),$
and one then optimizes over the choice of $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ to establish the lower bound.
3.2 Post-Interaction Reference and Two Complementary Approaches
In nonstationary settings, the reduction (7) typically breaks down: unlike the stationary case, there is no canonical separation condition that directly links regret to identifying the latent instance, since the underlying structure may vary over time. Motivated by the conditional bound (4), we introduce an interaction protocol with a post-interaction reference.
Protocol.
Draw an environment index $(u,b)\sim\mathrm{Unif}(\mathcal{V})$, where $\mathcal{V}:=\mathcal{U}\times\mathcal{B}$. Under the prior $\mathrm{Unif}(\mathcal{V})$, the random variables $u$ and $b$ are independent. Under the (unknown) environment $(u,b)$, let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ denote the interaction transcript, and let $P_{u,b}$ denote the law of $Z$ conditional on $(u,b)$. The key feature is that $b$ (the reference) is revealed to the estimator only after $Z$ has been generated. When $b$ is uninformative, this reduces to the classical formulation in Section 3.1.
Conditional versions of Fano’s lemma have been widely studied in statistics and information theory (Atia and Saligrama 2012, Scarlett and Cevher 2016, 2019), but have not, to our knowledge, been systematically developed as a general lower-bound mechanism for nonstationary bandits. We show that a post-interaction reference enables two complementary approaches for proving regret lower bounds.
Approach I: coordinate-wise identification error.
Analogous to (5), conditioning on $b$ yields
$I(u;Z\mid b)=\frac{1}{|\mathcal{B}|}\sum_{x\in\mathcal{B}}\frac{1}{|\mathcal{U}|}\sum_{\tau\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|\bar{P}_{x})\leq\frac{1}{|\mathcal{B}||\mathcal{U}|^{2}}\sum_{x\in\mathcal{B}}\sum_{\tau,\tau^{\prime}\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|P_{\tau^{\prime},x}),$
where $\bar{P}_{x}:=\frac{1}{|\mathcal{U}|}\sum_{\tau\in\mathcal{U}}P_{\tau,x}$. Plugging this into (4) gives
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{\frac{1}{|\mathcal{B}||\mathcal{U}|^{2}}\sum_{x\in\mathcal{B}}\sum_{\tau,\tau^{\prime}\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|P_{\tau^{\prime},x})+\log 2}{\log|\mathcal{U}|}.$ (8)
A particularly useful instantiation is when $\mathcal{V}$ has a product structure $\mathcal{V}=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{m}$. For any coordinate $i\in[m]$, we take
$\mathcal{U}=\mathcal{V}_{i},\qquad\mathcal{B}=\mathcal{V}_{-i}:=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{i-1}\times\mathcal{V}_{i+1}\times\cdots\times\mathcal{V}_{m}.$ (9)
This yields a coordinate-wise identification bound which can be paired with a lower bound on a partial regret component:
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\lesssim\mathbb{E}\big[\mathrm{Reg}_{T}^{(i)}(\pi,\nu)\big].$ (10)
Throughout Section 3, the notation $\lesssim$ may hide constants that depend on $T$ and $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. An intuitive interpretation of (8) and (10) is as follows: under an appropriate separation condition for a carefully constructed family of nonstationary instances, once the reference is revealed, any policy $\pi$ must incur a nontrivial coordinate-wise identification error. This nontrivial lower bound on the identification error can then be translated into a lower bound on the corresponding partial regret. If, moreover, the total regret decomposes as $\mathbb{E}[\mathrm{Reg}_{T}(\pi;\nu)]=\sum_{i=1}^{m}\mathbb{E}[\mathrm{Reg}_{T}^{(i)}(\pi;\nu)]$, then summing coordinate-wise bounds and optimizing over $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ yields Bayesian regret lower bounds. Since (10) is more general than (7)—both in allowing a broader separation condition and in relating identification to partial regret through the post-interaction reference—Approach I can be viewed as a strict extension of the interactive Fano methods for stationary bandits outlined in Section 3.1.
This coordinate-wise viewpoint is closely related to Assouad’s method for (passive) statistical estimation (Yu 1997). In the bandit setting, Chen et al. (2024) develop a general framework for stationary problems that recovers Assouad-type arguments. We show that the reference-aided coordinate-wise strategy extends naturally to nonstationary bandits: it yields the optimal dependence on the number of stationary segments $L$, and we leverage it to prove Theorem 1.
Approach II: coordinate-wise information cost .
We again adopt the product structure (9). Conditional on $(\mathcal{F}_{t-1},b)$, the action $a_{t}$ is drawn using the policy’s internal randomness, which is independent of $u$. Hence $I(u;a_{t}\mid\mathcal{F}_{t-1},b)=0$, and the chain rule yields
$I(u;Z\mid b)=\sum_{t=1}^{T}I\bigl(u;a_{t}\mid\mathcal{F}_{t-1},b\bigr)+\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr)=\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr).$
Consequently, (4) implies
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\!\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr)+\log 2}{\log|\mathcal{U}|}.$ (11)
The strength of (11) is that it yields regret lower bounds without requiring a direct implication of the form “(under an appropriate separation condition) small (partial) regret $\Rightarrow$ correct (coordinate-wise) identification,” an implication that is used both in the prior interactive Fano methods reviewed in Section 3.1 and in our Approach I, but that can fail in certain nonstationary constructions. Instead, we adopt an information cost perspective: we construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ and an estimator $\widehat{u}(Z,b)$ such that for some $\rho\in(0,1)$,
[…]
An intuitive interpretation of (11) and (12) is as follows: under any policy $\pi$, once the reference is revealed, the coordinate-wise identification error admits a nontrivial upper bound. Then, by (11), the learner must have acquired a sufficient amount of information about the underlying coordinate; in our construction, acquiring this information necessarily incurs some partial regret. In many nonstationary constructions, once $b$ is revealed, information about $u$ is acquired only through actions that incur regret, which makes (12) easier to verify than a direct reduction (10). We use this approach to prove Theorem 2.
We emphasize that, although classical KL change-of-measure arguments also link regret to information acquisition, they differ in an essential way from our approach. In the stationary setting, such arguments typically compare the laws of the interaction transcript under a small family of alternative environments and convert the resulting KL divergences into regret lower bounds (Bubeck et al. 2012, Garivier et al. 2019). Related ideas also appear in nonstationary bandits (Garivier and Moulines 2011). However, these arguments are usually organized around pairwise comparisons and are expressed directly in terms of KL divergence. By contrast, our lower-bound construction is genuinely multi-hypothesis, and it has a product structure across coordinates and is naturally formulated in terms of conditional mutual information. For this reason, a direct pairwise change-of-measure argument does not by itself provide the coordinate-wise decomposition and information-cost analysis required in our setting. We discuss this distinction further in the next section, where we explain why such classical arguments do not directly yield the desired logarithmic lower bounds for nonstationary satisficing bandits.
4 Regret Guarantees for Nonstationary Satisficing Bandits
In this section, we establish minimax satisficing regret guarantees for general nonstationary satisficing bandits. Sections 4.1 and 4.2 present the corresponding lower and upper bounds, respectively.
4.1 Lower Bounds
In the stationary setting, the minimax-optimal satisficing regret is constant under realizability (Michel et al. 2022, Feng et al. 2024, Rouyer et al. 2024). It is therefore natural to conjecture that, in a piecewise-stationary environment with $L$ segments, an $\mathcal{O}(L)$ guarantee might be achieved by restarting a stationary algorithm at each change-point (if the change-points can be efficiently detected). The main message of this subsection is that such an $\mathcal{O}(L)$ rate is not achievable in general when change-points are unknown: an additional $\log T$ factor is unavoidable.
A canonical hard class.
Let $\mathcal{E}_{L,\Delta}$ denote the class of two-armed, piecewise-stationary bandit instances with $L$ segments, where rewards are Gaussian with unit variance and the mean vector on each segment $\ell$ is either $(S+\Delta,S-\Delta)$ or $(S-\Delta,S+\Delta)$. In particular, the unique satisficing arm swaps at every change-point.
This class is canonical for our problem. It is the simplest symmetric realizable family that isolates the effect of nonstationarity: there are only two arms, the gap to the threshold is fixed at $\Delta$ on every segment, and each segment contains exactly one satisficing arm and one non-satisficing arm. Thus, the difficulty arises solely from the unknown change-points and the resulting switches in the identity of the satisficing arm. Furthermore, we emphasize that classical pairwise KL change-of-measure arguments do not directly yield the desired lower bound on this class. Indeed, if two instances in $\mathcal{E}_{L,\Delta}$ differ only in the location of one switch, then on the interval where the switch locations disagree. Hence, under unit-variance Gaussian rewards, each round in this disagreement interval contributes $2\Delta^{2}$ to the conditional KL divergence, regardless of which arm is pulled. As a result, the pairwise KL is driven primarily by the length of the disagreement interval, rather than by the number of regretful pulls. This breaks the usual mechanism behind classical KL-based lower bounds (Bubeck et al. 2012, Garivier et al. 2019, Garivier and Moulines 2011). We therefore use our post-interaction-reference framework to derive the sharp satisficing regret lower bound for $\mathcal{E}_{L,\Delta}$.
Theorem 1.
Suppose $L\geq 3$ and $\Delta^{2}T\geq L$. Then $\inf_{\pi\in\Pi}\sup_{\nu\in\mathcal{E}_{L,\Delta}}R_{T}^{S}(\pi;\nu)\gtrsim L\log(\Delta^{2}T/L)/\Delta.$
When $\Delta$ is a constant and $L\ll\Delta^{2}T$, Theorem 1 implies an $\Omega(L\log T)$ lower bound for general nonstationary satisficing bandits. Relative to the naive $\mathcal{O}(L)$ guess, the extra $\log T$ factor reflects the fact that the change-points are unknown: minimizing satisficing regret is at least as hard as localizing these change-points, which already incurs an $\Omega(L\log T)$ cost. The proof, deferred to Section 6.1, proceeds via Approach I in Section 3.2.
The proof of Theorem 1 is restricted to $L\geq 3$ because it relies on the particular instance construction (13), which is tailored to enforce the separation condition (14) and hence the reduction (10). When $L=2$ (i.e., the environment changes only once), it is unclear how to design an instance family admitting a comparable separation condition; see Figure 1. This leaves open whether the $\log T$ factor is also unavoidable when $L=2$. The next theorem closes this gap by establishing the same scaling for all $L\geq 2$, under a slightly stronger condition relating $\Delta$, $T$, and $L$.
separationseparationL=3L=3no separationL=2L=2
Figure 1: Schematic illustration: for $L=3$ one can create a separated local window to enable a clean identification reduction; for $L=2$ the analogous separation fails (motivating the information-budget approach).
Theorem 2.
Suppose $L\geq 2$ and $\Delta^{2}T\geq 13L$. Then $\inf_{\pi\in\Pi}\sup_{\nu\in\mathcal{E}_{L,\Delta}}R_{T}^{S}(\pi;\nu)\gtrsim L\log(\Delta^{2}T/L)/\Delta.$
Theorem 2 follows Approach II in Section 3.2; details are in Section 6.2. It also offers complementary intuition for the $\Omega(L\log T)$ scaling: regardless of the policy, one always acquires $\Omega(\log T)$ information per segment, since the reference can be used to drive the identification error below a constant. The information-cost bound then shows that obtaining this information necessarily incurs $\Omega(\log T)$ satisficing regret per segment.
4.2 Upper Bounds
We now present an algorithm whose satisficing regret matches the lower bounds above (up to problem-dependent constants). We begin by defining the windowed statistics used throughout.
Windowed statistics.
For any buffer $\mathcal{B}$ of length $m$, define the dyadic window set $\mathrm{WinSet}(\mathcal{B}):=\{2^{k}:2^{k}\leq m,\ k\in\mathbb{N}\}$. For each $w\in\mathrm{WinSet}(\mathcal{B})$, define the average of the most recent $w$ entries by $\mathrm{AvgLast}(\mathcal{B},w):=\frac{1}{w}\sum_{s=m-w+1}^{m}\mathcal{B}[s].$ Given a confidence radius $\beta(\cdot)$, define the windowed lower and upper confidence bounds $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B}):=\max_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)-\beta(w)\big\}$ and $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})=\min_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)+\beta(w)\big\}.$
Input: horizon $T$, threshold $S$, arms $[K]$, $\beta(w)\ \leftarrow\ \sqrt{2\big(4\log T+\log K\big)/w}$ for any $w\geq 1.$
[…]
for $t=1,2,\dots,T$ do
 if leader=None then
      Pull $a_{t}$ as the next arm in a fixed round–robin over $[K]$, observe $r_{t}$, append to $\mathcal{B}_{a_{t}}^{(\kappa_{a_{t}})}$.
[…]
 else
      Pull $a_{t}\leftarrow$ leader, observe $r_{t}$, append to $\mathcal{B}_{a_{t}}^{(\kappa_{a_{t}})}$.
[…]
Algorithm 1 Nonstationary–Sat (algorithm for nonstationary satisficing bandits)
Algorithm 1 maintains a leader state. When $\textsf{leader}=\texttt{None}$, it explores until it certifies that an arm is satisficing. When a leader is present, the algorithm exploits this arm while continuously testing whether it remains satisficing. The key difficulty in piecewise-stationary environments is that, after a change-point, stale samples can be misleading. The windowed bounds address this issue: $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})$ takes the most optimistic lower bound over dyadic suffix windows, while $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})$ takes the most pessimistic upper bound. Consequently, if an arm’s current mean exceeds $S$, then eventually $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})\geq S$; conversely, if the current leader becomes non-satisficing, then for a sufficiently recent window we have $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})<S$, triggering a reset and renewed exploration.
We now state the regret guarantee for Algorithm 1.
Theorem 3.
Under Assumption 1, let $\pi$ be the Algorithm 1. Then,
$R^{S}_{T}(\pi;\nu)\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}\big(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T$
where $\Delta_{a,\ell}:=(S-\mu_{a,\ell})_{+}$ and $\Delta_{\ell}^{*}:=\mu_{\ell}^{*}-S.$
Specializing to $\nu\in\mathcal{E}_{L,\Delta}$, Theorem 3 yields an $\mathcal{O}(L\log T/\Delta)$ upper bound, matching Theorems 1 and 2 up to $T$-independent factors. Moreover, the dependence $\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}$ also appears in stationary satisficing regret bounds; see, e.g., (Michel et al. 2022, Theorem 2).
The logarithmic dependence on $T$ arises from concentration: with $\beta(w)\asymp\sqrt{\log T/w}$, certifying $\mu>S$ (via $\mathrm{LCB}^{\operatorname{win}}\geq S$) or rejecting $\mu<S$ (via $\mathrm{UCB}^{\operatorname{win}}<S$) requires only $w=\mathcal{O}(\log T/\Delta^{2})$ samples, where $\Delta$ is the relevant gap to the threshold. The windowed bounds also prevent frequent resets: after a reset, if the environment remains unchanged, then with high probability a satisficing leader is not demoted and a non-satisficing candidate is not promoted. Thus, within each stationary segment, the algorithm spends only a logarithmic number of rounds (i) exploring until it finds a satisficing leader, or (ii) detecting that the current leader has fallen below $S$; on the remaining rounds it plays a satisficing arm and incurs no additional regret. The full proof of Theorem 3 is deferred to Section 6.3.
5 When Is Constant Satisficing Regret Achievable
It remains unclear which structural conditions permit constant satisficing regret in nonstationary $K$-armed bandits. In the stationary setting, constant satisficing regret is possible because the learner can identify satisficing and non-satisficing arms in finite expected time—independent of TT—by comparing empirical means to the threshold $S$; see (Michel et al. 2022, Appendix A).
A common feature of the hard instances behind the $\Omega(\log T)$ lower bounds in Section 4 is the presence of threshold crossings: an arm may be satisficing early on but later down-cross below $S$ at an unknown time, and the identity of the realizable arm may also switch across segments. Such changes force the learner to repeatedly certify satisficing arms, which incurs a logarithmic cost.
In this section, we identify a simple regime in which this obstacle disappears. Specifically, if (i) there exists an always-realizable arm that remains strictly satisficing throughout, and (ii) no arm ever transitions from satisficing to non-satisficing (i.e., no down-crossings), then constant satisficing regret—independent of both the horizon $T$ and the number of stationary segments LL—is achievable.
Formally, we impose the following assumption.
Assumption 2 (Always-Realizability).
$\exists a^{*}\in[K]$ such that $\mu_{a^{*},\ell}>S$ for all $\ell\in[0,L-1]$.
Assumption 3 (No Down-Crossing).
For each arm $a\in[K]$, the indicator sequence $\bigl\{\mathds{1}\{\mu_{t}(a)\geq S\}\bigr\}_{t\in[T]}$ is nondecreasing in $t$.
Under Assumptions 2 and 3, Simple-Sat (Algorithm 1 in Michel et al. (2022)), originally designed for the stationary setting, can be applied directly in the nonstationary case and achieves constant satisficing regret. Below, $\widehat{\mu}_{a}(t)$ denotes the empirical mean reward of arm $a$ based on observations collected up to time $t$.
Input: threshold $S$, arms $[K]$.
For time steps $t=1,\dots,K$, pull arm $a_{t}=t$.
for $t=K+1,K+2,\dots$ do
[…]
Algorithm 2 Simple-Sat (Michel et al. 2022)
Theorem 4.
Under Assumptions 2 and 3, let $\pi$ be Algorithm 2. Then,
$R_{T}^{S}(\pi;\nu)\lesssim\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\Big(1+\frac{1}{\Delta_{a,\min}^{2}}+\frac{1}{\Delta_{*}^{2}}\Big),$
where $\Delta_{a,\min}:=\min_{\ell\in[0,L-1]}\{\Delta_{a,\ell}:\ \Delta_{a,\ell}>0\}$, $\Delta_{a,\max}:=\max_{\ell\in[0,L-1]}\{\Delta_{a,\ell}:\ \Delta_{a,\ell}>0\}$ and $\Delta_{*}:=\min_{\ell\in[0,L-1]}(\mu_{a^{*},\ell}-S).$
Theorem 4 establishes a constant satisficing regret bound that does not depend on $T$ or $L$, and it recovers the stationary guarantee when $L=1$; see Michel et al. (2022). The key proof idea is that, under Assumptions 2 and 3, the always-realizable arm and any non-satisficing arm can be identified in finite expected time before the latter ever becomes satisficing. The full proof is deferred to Section 6.4.
We emphasize that Assumptions 2 and 3 are sufficient (but not necessary) for constant satisficing regret. Characterizing necessary and sufficient structural conditions remains an interesting direction for future work.
6 Proofs
6.1 Proof of Theorem 1
We prove Theorem 1 using Approach I in Section 3.2.
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
We assume Gaussian rewards of the form $r_{t}(a)=\mu_{t}(a)+\xi_{t}$, where $\xi_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume $L$ is odd and that 2T/(L1)2T/(L-1) is an integer. Define the block length m:=2T/(L1)m:=2T/(L-1) and partition the horizon into (L1)/2(L-1)/2 disjoint blocks $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\},$ $\forall\ell\in[(L-1)/2].$ Within each block $\ell$, there exists exactly one window of length $l$ whose location is unknown. Fix integers $n,l\geq 1$ with $nl\leq m-2$, and define the candidate window start times (within a block) by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw $\nu=(\nu_{1},\dots,\nu_{(L-1)/2}),$ where $\nu_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}).$ We interpret $\nu_{\ell}$ as the (within-block) start time of the special window in block $\ell$, and define the corresponding absolute start time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each time $t\in\mathcal{I}_{\ell}$, set
$(\mu_{t}(1),\mu_{t}(2))=\begin{cases}(S-\Delta,S+\Delta),&t\in[\widetilde{\nu}_{\ell},\widetilde{\nu}_{\ell}+l-1],\\ (S+\Delta,S-\Delta),&t\in\mathcal{I}_{\ell}\setminus[\widetilde{\nu}_{\ell},\widetilde{\nu}_{\ell}+l-1].\end{cases}$ (13)
This construction yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ be the complete interaction transcript, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{(L-1)/2})$ for the collection of all components except the $\ell$-th. A direct calculation shows that for any $\ell\in[(L-1)/2]$, any fixed $\tau_{-\ell}$, and any two distinct $\tau_{\ell}\neq\tau_{\ell}^{\prime}$, we have
$D_{\mathrm{KL}}(P_{\tau_{\ell},\tau_{-\ell}}\|P_{\tau^{\prime}_{\ell},\tau_{-\ell}})=4\Delta^{2}l.$ (14)
Therefore, by (8),
$\inf_{\widehat{\nu}_{\ell}(Z,\nu_{-\ell})}\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell}\bigr)$ $\geq 1-\frac{\frac{1}{n^{(L-3)/2}n^{2}}\sum_{\tau_{-\ell}}\sum_{\tau_{\ell},\tau^{\prime}_{\ell}}D_{\mathrm{KL}}(P_{\tau_{\ell},\tau_{-\ell}}\|P_{\tau^{\prime}_{\ell},\tau_{-\ell}})+\log 2}{\log n}$
$\geq 1-\frac{4\Delta^{2}l+\log 2}{\log n}\geq\frac{1}{2},$
where the last inequality holds by choosing $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$ and taking $n$ sufficiently large.
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$.
For each window $\tau\in\mathcal{T}$ in block $\ell$, construct the counter and the estimator:
$N_{2}^{(\ell)}(\tau):=\sum_{t=(\ell-1)m+\tau}^{(\ell-1)m+\tau+l-1}\mathds{1}\{a_{t}=2\},\qquad\widehat{\nu}_{\ell}(Z):=\arg\max_{\tau\in\mathcal{T}}N_{2}^{(\ell)}(\tau),$
with ties broken arbitrarily. Define the (blockwise) number of non-satisficing pulls: $M_{\ell}:=\sum_{t\in\mathcal{I}_{\ell}}\mathds{1}\{a_{t}\neq a_{t}^{*}\},$ where $a_{t}^{*}$ is the unique satisficing arm at time $t$. If $M_{\ell}\leq\frac{1}{4}l$, then inside the true window arm $2$ is satisficing for all $l$ rounds, so $N_{2}^{(\ell)}(\nu_{\ell})\ \geq\ l-M_{\ell}\ \geq\ \frac{3}{4}l.$ Consequently, $M_{\ell}\leq\frac{1}{4}l$ implies $\widehat{\nu}_{\ell}(Z)=\nu_{\ell}$, and thus
$\frac{1}{2}\leq\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z)\neq\nu_{\ell}\bigr)\leq\mathbb{P}\bigl(M_{\ell}>\tfrac{1}{4}l\bigr)$ $\leq\frac{4\mathbb{E}[M_{\ell}]}{l}\leq\frac{64\Delta^{2}\mathbb{E}[M_{\ell}]}{\log n}$
Then, $\mathbb{E}[M_{\ell}]\geq\frac{\log n}{128\Delta^{2}}$ for every $\ell\in[(L-1)/2]$. Since each pull of a non-satisficing arm incurs satisficing regret $\Delta$, $\mathbb{E}[R_{T}^{S}(\pi,\nu)]=\Delta\mathbb{E}[M]=\Delta\sum_{\ell\in[(L-1)/2]}\mathbb{E}[M_{\ell}]\geq\frac{(L-1)\log n}{256\Delta}.$
Optimize $n$.
Recall that $nl+2\leq 2T/(L-1)$ and $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$. Ignoring the ceiling, it suffices to choose $n\geq 2$ such that $n\log n\leq 16\Delta^{2}(\frac{2T}{L-1}-2).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}\geq L$, we further have $16\Delta^{2}(\frac{2T}{L-1}-2)>4.$ We will use the following lemma, whose proof is deferred to Section 6.1.1.
Lemma 1.
Let $y\geq 4$. Then there exists an integer $x\geq 2$ such that $x\log x\leq y$ and $\log x\geq\frac{1}{2}\log y.$
By Lemma 1, we finish the proof by choosing $n$ such that $\mathbb{E}[R_{T}^{S}(\pi,\nu)]\gtrsim\frac{L\log(\Delta^{2}T/L)}{\Delta}.$
6.1.1 Proof of Lemma 1
Let $x:=\lceil\sqrt{y}\rceil$. Then $x\geq\sqrt{y}$, so $\log x\geq\frac{1}{2}\log y$. It remains to show $x\log x\leq y$. Let $n:=\lfloor\sqrt{y}\rfloor$, so $n\geq 2$ and $y\in[n^{2},(n+1)^{2})$, which implies $x=\lceil\sqrt{y}\rceil\leq n+1$. Thus it suffices to prove
$(n+1)\log(n+1)\leq n^{2}\qquad\text{for all integers }n\geq 2.$ (15)
For $n=2$, (15) holds since $3\log 3\leq 4$. For $n\geq 3$, define the function $f(u):=u-1-\log(u+1)$ for $u\geq 3$. Then $f^{\prime}(u)=1-\frac{1}{u+1}=\frac{u}{u+1}>0,$ so $f$ is increasing on $3,\infty)$, and $f(n)\geq f(3)=2-\log 4>0$. Hence $\log(n+1)\leq n-1$ for all integers $n\geq 3$, and therefore $(n+1)\log(n+1)\leq(n+1)(n-1)=n^{2}-1\leq n^{2},$ proving (15). Since $y\geq n^{2}$, we conclude $x\log x\leq(n+1)\log(n+1)\leq n^{2}\leq y$. This completes the proof of Lemma 1.
6.2 Proof of Theorem 2
We prove Theorem 2 using Approach II in Section 3.2.
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
We assume Gaussian rewards of the form $r_{t}(a)=\mu_{t}(a)+\xi_{t}$, where $\xi_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume $L$ is even and that $2T/L$ is an integer. Define the block length $m:=2T/L$ and partition the horizon into $L/2$ disjoint blocks $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\},$ $\forall\ell\in[L/2].$ Within each block $\ell$, there is a single change-point $\nu_{\ell}$. Fix integers $n,l\geq 1$ with $nl\leq m-1$, and define the candidate change-point locations (within a block) by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw $\nu=(\nu_{1},\dots,\nu_{L/2}),$ where $\nu_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}).$ interpreting $\nu_{\ell}$ as the (within-block) change point in block $\ell$, and define the corresponding absolute time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each time $t\in\mathcal{I}_{\ell}$, set
$(\mu_{t}(1),\mu_{t}(2))=\begin{cases}(S-\Delta,S+\Delta),&t\in[(\ell-1)m+1,\widetilde{\nu}_{\ell}-1],\\ (S+\Delta,S-\Delta),&t\in[\widetilde{\nu}_{\ell},\ell m].\end{cases}$ (16)
This construction yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ be the complete interaction transcript, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{L/2})$ for the collection of all components except the $\ell$-th. Fix $\ell\in[L/2]$ and $t\in[T]$, and condition on $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})=(f,a,x)$. Define
$p:=\mathbb{P}(\widetilde{\nu}_{\ell}\leq t\mid f,a,x)\mathds{1}\{a=2\}+\mathbb{P}(\widetilde{\nu}_{\ell}>t\mid f,a,x)\mathds{1}\{a=1\},\quad P_{+}:=\mathcal{N}(S+\Delta,1),\quad P_{-}:=\mathcal{N}(S-\Delta,1).$
Given $(f,a,x)$, For any $t\in\mathcal{I}_{\ell}$, the conditional law of $r_{t}$ under the prior on $\nu_{\ell}$ is the mixture $Q:=(1-p)P_{+}+pP_{-}$. Moreover, conditionally on $(f,a,x)$, $r_{t}$ depends on $\nu_{\ell}$ only through the indicator $\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\}$. Thus, for any $t\in\mathcal{I}_{\ell}$,
$I(\nu_{\ell};r_{t}\mid F_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $=I\bigl(\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x\bigr)$
$=(1-p)D_{\mathrm{KL}}\bigl(P_{+}\|Q\bigr)+pD_{\mathrm{KL}}\bigl(P_{-}\|Q\bigr).$
If $t\notin\mathcal{I}_{\ell}$, then $r_{t}$ does not depend on $\nu_{\ell}$ given $(\mathcal{F}_{t-1},a_{t},b,u_{-\ell})$, and therefore
$I(\nu_{\ell};r_{t}\mid F_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)=0,\quad\forall t\in[T]\setminus\mathcal{I}_{\ell}.$
By convexity of $D_{\mathrm{KL}}$ in its second argument, $D_{\mathrm{KL}}(P_{+}\|Q)\leq pD_{\mathrm{KL}}(P_{+}\|P_{-})$ and $D_{\mathrm{KL}}(P_{-}\|Q)\leq(1-p)D_{\mathrm{KL}}(P_{-}\|P_{+}).$ Thus, for any $t\in\mathcal{I}_{\ell}$,
$I(\nu_{\ell};r_{t}\mid F_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $\leq(1-p)pD_{\mathrm{KL}}(P_{+}\|P_{-})+p(1-p)D_{\mathrm{KL}}(P_{-}\|P_{+})$
$\leq p(D_{\mathrm{KL}}(P_{+}\|P_{-})+D_{\mathrm{KL}}(P_{-}\|P_{+}))=4\Delta^{2}p.$
For any $t\in\mathcal{I}_{\ell}$, taking expectation over $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$ yields
$I(\nu_{\ell};r_{t}\mid F_{t-1},a_{t},\nu_{-\ell})$ $\leq 4\Delta^{2}\mathbb{E}\Bigl[\mathbb{P}(\widetilde{\nu}_{\ell}\leq t\mid F_{t-1},a_{t},\nu_{-\ell})\mathds{1}\{a_{t}=2\}+\mathbb{P}(\widetilde{\nu}_{\ell}>t\mid F_{t-1},a_{t},\nu_{-\ell})\mathds{1}\{a_{t}=1\}\Bigr]$
$=4\Delta^{2}\mathbb{E}\Bigl[\mathds{1}\{\widetilde{\nu}_{\ell}\leq t,a_{t}=2\}+\mathds{1}\{\widetilde{\nu}_{\ell}>t,a_{t}=1\}\Bigr].$
Finally, for any $\ell\in[L/2]$,
$I(u_{\ell};Z\mid B,u_{-\ell})\leq 4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr].$ (17)
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$.
Fix $\ell\in[L/2]$. For each candidate change-point $\tau\in\mathcal{T}$, define
$B_{\ell,\tau}:=\bigl[(\ell-1)m+\tau,\;(\ell-1)m+\tau+l-1\bigr],\qquad N_{\ell,\tau}(Z):=\sum_{t\in B_{\ell,\tau}}\mathds{1}\{a_{t}=1\}.$
Fix an integer $r$ with $1\leq r\leq l$. For each $\tau$ such that $N_{\ell,\tau}(Z)\geq r$, let $\overline{X}_{\ell,\tau}(r)$ denote the average of the first $r$ rewards of arm $1$ observed within $B_{\ell,\tau}$. For each $\tau$ such that $N_{\ell,\tau}(Z)<r$, let $\overline{X}^{\prime}_{\ell,\tau}(l-r)$ denote the average of the first lrl-r rewards of arm $2$ observed within $B_{\ell,\tau}$. Define the estimators
$\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)\geq r\ \text{ and }\ \overline{X}_{\ell,\tau}(r)\geq S\Bigr\}$
$\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)<r\ \text{ and }\ \overline{X}_{\ell,\tau}^{\prime}(l-r)\leq S\Bigr\}$
We next state a lemma that bounds the error probabilities of the two estimators above; its proof is deferred to Section 6.2.1.
Lemma 2.
Fix integers $l,r\in\mathbb{N}$ with $1\leq r\leq l$, and $\ell\in[L/2]$. Then,
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}(N_{\ell,\nu_{\ell}}(Z)<r)+n\exp(-r\Delta^{2}/2),$
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}\big(N_{\ell,u_{\ell}}(Z)\geq r\big)+n\exp(-(l-r)\Delta^{2}/2).$
Construct the final estimator by flipping an independent coin $\Xi\sim\mathrm{Unif}\{0,1\}$. If $\Xi=0$, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$; otherwise, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$. Combining Lemma 2 with the choices $r=\lceil l/2\rceil$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$, we obtain
$\mathbb{P}(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell})\ \leq 1/2+n\exp(-(l-1)\Delta^{2}/2)\ \leq 1/2+1/n^{2}\leq 3/4,$ (18)
Combining (17) and (18), we obtain
$4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{1}{4}\log n-\log 2,\quad\forall\ell\in[L/2].$
Consequently,
$\mathbb{E}[R_{T}^{S}(\pi;\nu)]=\Delta\cdot\sum_{\ell\in[L/2]}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{L}{8\Delta}\Bigl(\frac{1}{4}\log n-\log 2\Bigr).$
Optimize $n$.
Recall the constraints $nl\leq 2T/L-1$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$. Ignoring the ceiling for simplicity, it suffices to choose $n\geq 2$ such that $n\log n\;\leq\;\frac{\Delta^{2}}{6}(\frac{2T}{L}-1).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}/L\geq 13$, we further have $\frac{\Delta^{2}}{6}(\frac{2T}{L}-1)\geq 13/3-1/32>4.$ By Lemma 1, we may choose $n\geq 2$ so that $\mathbb{E}\!\left[R_{T}^{S}(\pi;\nu)\right]\gtrsim\frac{L}{\Delta}\log(\Delta^{2}T/L)$. This completes the proof of Theorem 2.
6.2.1 Proof of Lemma 2
In this section, we prove the upper bound on $\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ in Lemma 2; the second bound follows by an analogous argument. Let $E:=\{N_{\ell,\nu_{\ell}}(Z)<r\}$. Then
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})\leq\mathbb{P}(E)+\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}}).$
On the event $E^{\mathrm{c}}$, the estimator $\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ can make error only if either
  • the block at the true change-point is not declared, i.e., $\overline{X}_{\ell,\nu_{\ell}}<S$; or
  • some earlier candidate is declared first, i.e., $\overline{X}_{\ell,\tau}\geq S$ for some $\tau<\nu_{\ell}$ with $N_{\ell,\tau}(Z)\geq r$.
For the first event, since $\overline{X}_{\ell,\nu_{\ell}}\sim\mathcal{N}(S+\Delta,1/r)$ on $E^{\mathrm{c}}$, the Gaussian tail bound gives
$\mathbb{P}(\overline{X}_{\ell,\nu_{\ell}}<S)=\mathbb{P}\left(\overline{X}_{\ell,\nu_{\ell}}-(S+\Delta)<-\Delta\right)\leq\exp(-r\Delta^{2}/2).$
Next, fix any $\tau<\nu_{\ell}$. On $E^{\mathrm{c}}\cap\{N_{\ell,\tau}(Z)\geq r\}$, we have $\overline{X}_{\ell,\tau}\sim\mathcal{N}(S-\Delta,1/r)$, and hence
$\mathbb{P}(\overline{X}_{\ell,\tau}\geq S)=\mathbb{P}(\overline{X}_{\ell,\tau}-(S-\Delta)\geq\Delta)\leq\exp(-r\Delta^{2}/2).$
Taking a union bound over the at most n1n-1 indices $\tau\in\mathcal{T}$ with $\tau<\nu_{\ell}$, we obtain
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}})\leq(n-1)\exp(-r\Delta^{2}/2)+\exp(-r\Delta^{2}/2)=n\exp(-r\Delta^{2}/2).$
6.3 Proof of Theorem 3
By the definition of the satisficing regret in (2), we have
$R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a=1}^{K}\Delta_{a,\ell}\mathbb{P}_{\pi,\nu}(a_{t}=a)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a),$ (19)
where, in the last equality, we simplify the notation by omitting the dependence on the policy $\pi$ and the environment $\nu$. We first introduce some useful notation. For $\Delta>0$, define
$n_{\Delta}:=\min\Bigl\{n\in[T]:\sqrt{\frac{2(4\log T+\log K)}{n}}\leq\Delta/2\Bigr\}\leq c_{1}\frac{\log TK}{\Delta^{2}},$
where $c_{1}>0$ is a universal constant. Consequently, we have $\beta(n)\leq\Delta/2$ for all $n$ such that $n_{\Delta}\leq n\leq T$. Define the event $\mathcal{E}_{\operatorname{win}}$ that for every arm $a$, epoch $\kappa$, time $t\leq T$, and every dyadic window $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$ whose last $w$ samples lie entirely inside a stationary segment $\ell$, we have $\Big|\mathrm{AvgLast}\big(\mathcal{B}_{a}^{(\kappa)}(t),w\big)-\mu_{a,\ell}\Big|\ \leq\ \beta(w).$ By Hoeffding’s inequality for $1$-sub-Gaussian variables, for fixed $(a,\kappa,t,w)$ whose last $w$ samples lie in some stationary segment $\ell$,
$\mathbb{P}\left(\Big|\mathrm{AvgLast}\big(\mathcal{B}_{a}^{(\kappa)}(t),w\big)-\mu_{a,\ell}\Big|\geq\beta(w)\right)$ $\leq 2\exp\big(-\frac{w}{2}\beta(w)^{2}\big)=2\exp(-4\log T-\log K)=\frac{2}{KT^{4}}.$
Taking a union bound over all $a\in[K]$, epochs $\kappa\leq T$, times $t\leq T$, and windows $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$, and using that $|\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))|\leq T$, we obtain $\mathcal{E}_{\operatorname{win}}^{c}\leq\frac{2}{T}.$ On the event $\mathcal{E}_{\operatorname{win}}$, the windowed confidence bounds behave as follows:
  • Fast rejection. If $\mu_{a,\ell}=S-\Delta$ and $\mathcal{B}_{a}^{(\kappa)}$ contains a suffix window of length $w\geq 2n_{\Delta}$ entirely within segment $\ell$, then $\mathrm{UCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\leq S$. Hence any leader whose mean stays $\leq S-\Delta$ in an epoch is demoted after at most $2n_{\Delta}$ pulls within a single segment.
  • Fast certification. Similarly, any arm whose mean stays $\geq S+\Delta$ in an epoch is certified as satisficing once it accumulates $2n_{\Delta}$ pulls within a single segment.
  • At most one false leader per segment. Within a fixed segment $\ell$, any non-satisficing arm can be certified as leader at most once: after it is demoted its buffer resets, and while its mean remains $\leq S-\Delta$, it cannot satisfy $\mathrm{LCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\geq S$ again before the segment ends.
  • No demotion of the optimal arm. Within segment $\ell$, once an optimal (satisficing) arm is certified as leader, it is never demoted before the segment ends.
On the event $\mathcal{E}_{\operatorname{win}}$, the exploration length on each segment $\ell$ is at most $2Kn_{\Delta_{\ell}^{*}}$: during exploration the algorithm runs a fixed round–robin schedule, and once the optimal arm on segment $\ell$ has been pulled $2n_{\Delta_{\ell}^{*}}$ times, it is certified as satisficing and exploration terminates. Moreover, since every non-satisficing arm $a$ can be certified as leader at most once in segment $\ell$ and, once certified, is demoted after at most $2n_{\Delta_{a,\ell}}$ pulls, the regret incurred on segment $\ell$ is bounded by
$\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log KT\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T$
up to a universal constant. Therefore, we have
$R_{T}^{S}(\pi;\nu)$ $=\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\mid\mathcal{E}_{\operatorname{win}}\Big]\mathbb{P}(\mathcal{E}_{\operatorname{win}})+\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\mid\mathcal{E}_{\operatorname{win}}^{c}\Big]\mathbb{P}(\mathcal{E}_{\operatorname{win}}^{c})$
$\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T+T\cdot\frac{2}{T}\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T,$
thereby completing the proof of Theorem 3.
6.4 Proof of Theorem 4
In this section, we prove Theorem 4 by adapting the proof of Theorem 1 in Michel et al. (2022) from the stationary setting to our nonstationary setting under Assumptions 2 and 3. We work with the same regret representation as in equation (19): $R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a).$ Following the event decomposition used in Appendix A of Michel et al. (2022), we write, for each arm $a$ and time $t$, $\left\{a_{t}=a\right\}\subset\{t=a\}\cup\left\{a_{t}=a,Z_{t}^{c}\right\}\cup\left\{a_{t}=a,Z_{t}\right\},$ where $Z_{t}:=\left\{\forall a\in[K],\hat{\mu}_{t}(a)<S\right\}$. For the first event, we have
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(t=a)\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}.$
For each arm $a\in[K]\setminus a^{*}$, define $T(a):=\max\{t\in[T]:\mu_{t}(a)<S\}$. For $n\leq T(a)$, let $\hat{\mu}_{a,n}$ denote the empirical mean of arm $a$ based on the first $n$ pulls of $a$ that occur strictly before time $T(a)$. For the second event, Assumption 3 implies that
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t}^{c})$ $\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\mathbb{P}(\hat{\mu}_{a,n}(a)\geq S)$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\exp(-\frac{n\Delta_{a,\min}^{2}}{2})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{a,\min}^{2}},$
Define $\hat{\mu}_{a^{*},n}$ denote the empirical mean of arm $a^{*}$ based on the first $n$ pulls of $a^{*}$. For the last event, Assumption 2 implies that
$\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t})$ $=\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a^{*},Z_{t})$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T}\mathbb{P}(\hat{\mu}_{a^{*},n}\leq S)\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{*}^{2}}.$
Combining three events together, we complete the proof of Theorem 4.
7 Conclusion
This paper studies satisficing regret in nonstationary bandits. Our results reveal a sharp contrast with the stationary realizable setting: once the reward process is allowed to change at unknown times, constant satisficing regret is no longer achievable in general. In particular, in the realizable piecewise-stationary setting with $L$ stationary segments, we show that the optimal regret scales as $\Theta(L\log T)$ whenever $L\geq 2$.
Our lower bounds are based on a new post-interaction-reference framework, which strictly extends interactive Fano methods from stationary bandits to nonstationary settings.
We also show that constant satisficing regret can still be recovered under additional structural assumptions. An important direction for future work is to characterize more precisely which structural conditions are sufficient, and which are necessary, for constant satisficing regret under nonstationarity.
[…]
Q. Xie and Y. Zhang are supported in part by National Science Foundation (NSF) grants CNS-1955997 and EPCN-2339794 and EPCN-2432546.