### 25 Apr 2026 v2

# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

[…]
On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization
[…]
On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization
[…]
Motivated by the principle of satisficing in decision-making, we study satisficing regret guarantees for nonstationary $K$-armed bandits. We show that in the general realizable, piecewise-stationary setting with $L$ stationary segments, the optimal regret is $\Theta(L\log T)$ as long as $L\geq 2$. This stands in sharp contrast to the case of $L=1$ (i.e., the stationary setting), where a $T$-independent $\Theta(1)$ satisficing regret is achievable under realizability. In other words, the optimal regret has to scale with $T$ even if just a little nonstationarity presents. A key ingredient in our analysis is a novel Fano-based framework tailored to nonstationary bandits via a post-interaction reference construction. This framework strictly extends the classical Fano method for passive estimation as well as recent interactive Fano techniques for stationary bandits. As a complement, we also discuss a special regime in which constant satisficing regret is again possible.
[…]
nonstationary bandits, satisficing regret, fano methods, post-interaction reference
1 Introduction
The multi-armed bandit (MAB) problem is a classical framework for studying the exploration–exploitation trade-off in sequential decision-making. In the standard MAB setting, a learner repeatedly chooses among $K$ arms with unknown, arm-specific stationary reward distributions, with the goal of minimizing cumulative regret (Robbins 1952, Lai and Robbins 1985). For the canonical MAB problem and its many variants, a rich literature has established sharp performance guarantees; see, for example, Lattimore and Szepesvári (2020).
While identifying the optimal action is essential in some high-stakes settings, many practical applications only require a “good-enough,” or satisficing, decision. Introduced by Simon (1956), satisficing aims for an acceptable solution rather than the absolute best one. This perspective is common in practice. For instance, recommendation and advertising systems often deploy a candidate only if its KPI exceeds a threshold $S$ with sufficient confidence (Kohavi et al. 2020, Xu et al. 2015), and inventory control typically targets a desired service level (e.g., 90–95%) to mitigate stockout costs (Fitzsimons 2000, Jing and Lewis 2011).
These considerations motivate the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2022, Rouyer et al. 2024, Feng et al. 2024, Hajiabolhassan and Ortner 2025), where the learner aims to pull a satisficing arm—an arm whose mean reward is no less than a prescribed threshold SS—as often as possible. Performance is measured by the satisficing regret, defined as the cumulative shortfall of the selected arm’s mean reward relative to $S$. To the best of our knowledge, existing results on satisficing regret primarily focus on stationary environments: Tamatsukuri and Takahashi (2019), Michel et al. (2022), Rouyer et al. (2024) study the $K$-armed bandit setting, Feng et al. (2024) extend the framework to more general bandit optimization problems, and Hajiabolhassan and Ortner (2025) study the average-reward Markov decision process setting. Under a realizability condition—namely, that at least one (initially) unknown arm is strictly satisficing with mean exceeding SS—these works show that the satisficing regret can be bounded by a constant independent of the decision horizon $T$.
In many real-world scenarios, however, stationarity fails: consumer preferences drift with seasons and trends (Ardeshiri et al. 2019), and platform marketplaces (e.g., ride-hailing) experience demand surges and changing competitive conditions (Battifarano and Qian 2019). As a result, an arm’s reward distribution may evolve over time. Despite the growing literature on satisficing regret in stationary environments, much less is understood in the nonstationary setting that frequently arises in practice. Motivated by this gap, we study satisficing regret minimization under nonstationarity, focusing on piecewise-stationary bandits with $L$ stationary segments. Building on the stationary case (corresponding to $L=1$), it is natural to ask:
Can we similarly achieve a $T$-independent satisficing regret in the piecewise-stationary setting?
At first glance, this seems plausible: if the change points were known in advance, one could simply restart a stationary satisficing algorithm on each segment, leading to an $\mathcal{O}(L)$ satisficing regret.
1.1 Our Contributions
In this work, however, we provide a negative answer to the above conjecture. Specifically,
  • In Section 4, we establish an $\Omega\bigl(L\log(T/L)\bigr)$ lower bound on the satisficing regret for nonstationary $K$-armed bandits. We then propose Algorithm 1, which achieves an almost matching upper bound of $\mathcal{O}(L\log T)$. Together, these results show that the additional $T$-dependent $\log T$ factor is unavoidable in general when the change points are unknown, thereby ruling out the conjectured $\mathcal{O}(L)$ satisficing regret guarantee.
  • To establish the lower bounds above, we extend recent interactive Fano techniques for bandit regret lower bounds (Chen et al. 2024, Shufaro et al. 2024) from stationary to nonstationary environments. More importantly, we introduce a post-interaction reference framework that strictly generalizes the classical Fano approach to make it well suited to nonstationary bandits. In this framework, a key distinction compared to the Fano techniques is that part of the (initially) unknown environment parameters (the “reference”) are revealed to the estimator after the entire horizon of length $T$. This leads to two novel complementary paths in lower bounding the regret (Section 3.2): (i) reducing a portion of the regret to an identification error under a suitable separation condition (used in establishing the lower bound when $L\geq 3$); and (ii) interpreting a portion of the regret as an information cost (used for $L=2$). We also note that this framework may extend to broader classes of nonstationary bandit problems and may be of independent interest.
  • In Section 5, we complement the general negative results by identifying a special case in which constant satisficing regret is achievable. Specifically, when there exists an always-realizable arm and no satisficing arm ever becomes non-satisficing, following the results in Michel et al. (2022) can yield a constant satisficing regret that is independent of both $T$ and $L$.
1.2 Related Work
Nonstationary Bandits. Most work on nonstationary bandits evaluates performance through dynamic regret, which compares the learner with the best arm at each round. For piecewise-stationary environments with at most $L$ segments, sliding-window and change-point-based methods achieve $\widetilde{\mathcal{O}}(\sqrt{LT})$ dynamic regret (Garivier and Moulines 2011, Auer et al. 2019). Under a variation-budget model with total drift $D$, the optimal rate is $\widetilde{\mathcal{O}}(D^{1/3}T^{2/3})$ (Besbes et al. 2014, Cheung et al. 2019), and Wei and Luo (2021) obtain the unified guarantee $\widetilde{\mathcal{O}}(\min\{\sqrt{LT},D^{1/3}T^{2/3}\})$. These results address a stronger benchmark than ours: dynamic regret competes with the instantaneous best arm, whereas our objective only requires selecting an arm whose mean exceeds a fixed threshold.
Stationary Satisficing Regret. Our work is most closely related to the literature on threshold-based satisficing regret in stationary bandits. Michel et al. (2022), Rouyer et al. (2024), and Feng et al. (2024) study cumulative satisficing regret under a fixed threshold and show that, under realizability, constant regret is achievable in stationary settings. Our paper shows that this phenomenon changes sharply under nonstationarity: once the environment may change at unknown times, constant satisficing regret is no longer achievable in general.
Other Satisficing Notions. In stationary settings, threshold-based objectives have also been studied in pure exploration under the name thresholding bandits (Locatelli et al. 2016, Mukherjee et al. 2017). There, the goal is to identify arms above a threshold, with performance measured by error probability or sample complexity rather than cumulative satisficing regret. Moreover, a naive explore–then–exploit reduction does not yield optimal cumulative satisficing regret even in the stationary case (Michel et al. 2022, Feng et al. 2024).
Russo and Van Roy (2018) introduce a discounted notion of satisficing in stationary bandits, motivated by time-sensitive learning. Their setting and benchmark differ from ours: they study discounted regret rather than cumulative threshold shortfall.
For nonstationary bandits, as highlighted in Liu et al. (2023b, a), the standard notion of dynamic regret can steer algorithm design toward overly aggressive exploration; it can therefore be more meaningful to compete against a satisficing benchmark. Min and Russo (2023) study nonstationary bandits in a Bayesian model and introduce a notion of satisficing based on competing with a satisficing action sequence $A^{\dagger}$ (Min and Russo 2023, Sec. 7). Their “satisficing regret” is defined as regret relative to $A^{\dagger}$. In contrast, we study a threshold-based notion of satisficing regret, defined as the cumulative shortfall from a fixed threshold $S$. Our formulation does not require that any arm achieves mean exactly $S$ at each time. Even if one sets $A^{\dagger}$ to be a realizable arm sequence, the two regret notions are not directly comparable: regret relative to $A^{\dagger}$ can be negative at times when the learner selects an arm with strictly larger mean than $A_{t}^{\dagger}$, whereas our satisficing regret is nonnegative by definition. Finally, Min and Russo (2023) establishes asymptotic lower bounds for the classical dynamic regret, while we provide minimax-optimal finite-time guarantees for threshold-based satisficing regret.
1.3 Notation
For integers $M\geq 1$, define $M]:=\{1,2,\ldots,M\}$. For integers $0\leq N\leq M$, define $N,M]:=\{N,N+1,\ldots,M\}$. For a buffer $\mathcal{B}=(b_{1},\ldots,b_{m})$ of length $m$, we write $\mathcal{B}[s]:=b_{s}$ for its $s$-th entry, $s\in[m]$. For a set $\mathcal{A}$, we use $|\mathcal{A}|$ to denote its cardinality, and $\mathrm{Unif}(\mathcal{A})$ for the uniform distribution on $\mathcal{A}$. For random variables $(X,Y)$ with joint law $P_{X,Y}$, the mutual information is
$I(X;Y):=\mathbb{E}_{X}\left[D_{\mathrm{KL}}\bigl(P_{Y\mid X}\,\|\,P_{Y}\bigr)\right],$ (1)
[…]
2 Problem Formulation
We study a nonstationary $K$-armed bandit over a horizon $T\in\mathbb{N}$, with $K\lesssim T$. At each round $t\in[T]$, each arm $a\in[K]$ has an unknown mean reward $\mu_{t}(a)\in[0,1]$. For integers $1\leq s\leq t\leq T$, define the number of stationary segments on $[s,t]$ as
$L[s,t]:=1+\sum_{j=s}^{t-1}\mathds{1}\left\{\exists a\in[K]\text{ such that }\ \mu_{j}(a)\neq\mu_{j+1}(a)\right\}.$
Let $L:=L[1,T]$ denote the total number of stationary segments and $1=T_{0}<T_{1}<\cdots<T_{L}=T+1$ be the associated change-points that partition the horizon into $L$ stationary segments, where segment $\ell\in\{0,\dots,L-1\}$ is the interval $T_{\ell},\,T_{\ell+1}-1]$. A (possibly randomized) policy $\pi$ selects an action $a_{t}\in[K]$ at time $t$, adapted to the natural filtration $\mathcal{F}_{t-1}:=\sigma(a_{1},r_{1},\dots,a_{t-1},r_{t-1})$. Let $\Pi$ denote the class of all such policies. Conditional on $a_{t}$, the learner observes a reward $r_{t}=\mu_{t}(a_{t})+\xi_{t},$ where $(\xi_{t})_{t\geq 1}$ are independent, zero-mean, $1$-sub-Gaussian noises. For each segment $\ell\in\{0,\dots,L-1\}$, let $\mu_{a,\ell}:=\mu_{T_{\ell}}(a)$ denote the mean reward of arm $a\in[K]$ on segment $\ell$, and define the segment-wise optimal mean $\mu_{\ell}^{*}:=\max_{a\in[K]}\mu_{a,\ell}$.
The decision maker specifies a threshold $S\in\mathbb{R}$, and performance is judged by whether the chosen arm meets this level. An arm $a$ is satisficing at time $t$ if $\mu_{t}(a)\geq S$. The one-step satisficing regret is $(S-\mu_{t}(a_{t}))_{+}:=\max\{0,S-\mu_{t}(a_{t})\}$, and the cumulative satisficing regret under policy $\pi$ is
$R_{T}^{S}(\pi;\nu):=\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\Big],$ (2)
where $\nu$ denotes the underlying nonstationary environment and $\mathbb{E}_{\pi,\nu}$ is over the policy’s randomness and the observation noise. Throughout, we focus on the standard realizable case as specified below (Michel et al. 2022, Feng et al. 2024).
Assumption 1 (Realizability).
For every segment $\ell\in\{0,\dots,L-1\}$, there exists at least one arm with mean strictly above the threshold, i.e., $\mu_{\ell}^{*}>S$.
3 A Fano-Based Framework for Nonstationary Bandit Problems
In this section, we detail our new technical development based on Fano’s lemma (Reza 1994, Theorem 2.10) . We start by reviewing its usage for lower bounds in stationary bandits (Section 3.1), and then derive the post-interaction reference framework (Section 3.2). This yields two powerful approaches that are particularly suitable for deriving regret lower bounds for nonstationary bandits.
Fano’s lemma is a classical tool for establishing minimax lower bounds in multi-hypothesis testing. Let $\mathcal{V}$ be a finite index set, $\nu\sim\mathrm{Unif}(\mathcal{V})$ be a latent instance index, and $Z$ be the data generated by an algorithm interacting with the instance indexed by $\nu$, Fano’s lemma lower bounds the probability of being unable to correctly identify $\nu$ with $Z$ in terms of the mutual information $I(\nu;Z)$:
$\inf_{\widehat{\nu}(Z)}\mathbb{P}\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\geq 1-\frac{I(\nu;Z)+\log 2}{\log|\mathcal{V}|}.$ (3)
Being initially designed for multi-hypothesis testing, recent works (Chen et al. 2024, Shufaro et al. 2024) build on it to develop a general framework for proving regret lower bounds in stationary bandit problems.
Conditional Fano. Aside from the above, it is often useful to apply Fano’s lemma in a structured setting where the index takes the form $\nu=(u,b)\in\mathcal{U}\times\mathcal{B}$. Here, $b$ is a “reference” variable independent of $u$. The observation $Z$ may depend on $b$, and $b$ is revealed to the estimator before it outputs $\widehat{u}$. A conditional form of Fano’s inequality (Scarlett and Cevher 2019, Theorem 3) then follows directly from (3):
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{I(u;Z\mid b)+\log 2}{\log|\mathcal{U}|}.$ (4)
3.1 Classical Fano Reduction: Small Regret Implies Correct Identification
For a bandit problem, we write out $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ as the full interaction transcript obtained by running $\pi$ for $T$ rounds under the environment indexed by $\nu$. Let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$ and $\bar{P}:=\frac{1}{|\mathcal{V}|}(\sum_{\tau\in\mathcal{V}}P_{\tau})$ be the resulting mixture distribution. By definition of mutual information and the convexity of $D_{\mathrm{KL}}(P\|\cdot)$ in its second argument (Cover 1999), we have
$I(\nu;Z)=\frac{1}{|\mathcal{V}|}\sum_{\tau\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|\bar{P}\right)\leq\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right).$ (5)
Combining (5) with (3) yields
$\inf_{\widehat{\nu}(Z)}\mathbb{P}\!\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\geq 1-\frac{\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right)+\log 2}{\log|\mathcal{V}|}.$ (6)
In stationary bandits, a standard route to regret lower bounds is to reduce small regret to accurate identification of the underlying instance $\nu$. Specifically, one constructs an instance family $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ satisfying a separation condition (among different arms’ mean rewards) at level $\Delta$ (see, e.g., Chen et al. 2024, Eq. (10)), which enables the construction of an estimator $\widehat{\nu}(Z)$ from the transcript $Z$ such that
$\mathbb{P}\bigl(\widehat{\nu}(Z)\neq\nu\bigr)\leq\mathbb{P}\bigl(\mathrm{Reg}_{T}(\pi,\nu)\geq T\Delta\bigl),$ (7)
where $\mathrm{Reg}_{T}$ denotes the regret of interest. By Markov’s inequality and (6), the regret under the uniform prior satisfies
$\mathbb{E}\!\left[\mathrm{Reg}_{T}(\pi,\nu)\right]\;\geq\;T\Delta\cdot\Bigl(1-\frac{\frac{1}{|\mathcal{V}|^{2}}\sum_{\tau,\tau^{\prime}\in\mathcal{V}}D_{\mathrm{KL}}\!\left(P_{\tau}\middle\|P_{\tau^{\prime}}\right)+\log 2}{\log|\mathcal{V}|}\Bigr),$
and one then optimizes over the choice of $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ to establish the lower bound.
3.2 Post-Interaction Reference and Two Complementary Approaches
In nonstationary bandits, however, the reduction (7) typically fails because the canonical separation condition may not hold due to changes in the arms’ mean rewards. To overcome this, we introduce the post-interaction reference based on the conditional bound (4). To the best of our knowledge, our framework is the first systematic application of the conditional Fano’s method to establish lower bounds in (nonstationary) bandits.
Post-Interaction Reference Protocol. Draw an environment index $(u,b)\sim\mathrm{Unif}(\mathcal{V})$, where $\mathcal{V}:=\mathcal{U}\times\mathcal{B}$. Under the prior $\mathrm{Unif}(\mathcal{V})$, the random variables $u$ and $b$ are independent. Under the (unknown) environment $(u,b)$, we let $P_{u,b}$ denote the law of $Z$ conditional on $(u,b)$. The key feature of the post-interaction reference framework is that $b$ (the reference) is revealed to the estimator only after $Z$ has been generated. This enables the below two complementary approaches for proving regret lower bounds.
Approach I: Coordinate-Wise Identification Error. Analogous to (5), conditioning on $b$ yields
$I(u;Z\mid b)=\frac{1}{|\mathcal{B}|}\sum_{x\in\mathcal{B}}\frac{1}{|\mathcal{U}|}\sum_{\tau\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|\bar{P}_{x})\leq\frac{1}{|\mathcal{B}||\mathcal{U}|^{2}}\sum_{x\in\mathcal{B}}\sum_{\tau,\tau^{\prime}\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|P_{\tau^{\prime},x}),$
where $\bar{P}_{x}:=\frac{1}{|\mathcal{U}|}\sum_{\tau\in\mathcal{U}}P_{\tau,x}$. Plugging this into (4) gives
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{\frac{1}{|\mathcal{B}||\mathcal{U}|^{2}}\sum_{x\in\mathcal{B}}\sum_{\tau,\tau^{\prime}\in\mathcal{U}}D_{\mathrm{KL}}(P_{\tau,x}\|P_{\tau^{\prime},x})+\log 2}{\log|\mathcal{U}|}.$ (8)
A particularly useful instantiation is when $\mathcal{V}$ has a product structure $\mathcal{V}=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{m}$. For any coordinate $i\in[m]$, we can take
$\mathcal{U}=\mathcal{V}_{i},\qquad\mathcal{B}=\mathcal{V}_{-i}:=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{i-1}\times\mathcal{V}_{i+1}\times\cdots\times\mathcal{V}_{m}.$ (9)
This coordinate-wise construction is particularly well suited to nonstationary bandits: each factor $\mathcal{V}_{i}$ can be viewed, roughly, as an instance family associated with one stationary segment. This structure is key to obtaining the optimal dependence on the number of stationary segments. In particular, it yields a coordinate-wise identification bound that can be paired with a lower bound on a corresponding partial regret term:
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\lesssim\mathbb{E}\big[\mathrm{Reg}_{T}^{(i)}(\pi,\nu)\big],$ (10)
and the total regret decomposes as $\mathbb{E}[\mathrm{Reg}_{T}(\pi;\nu)]=\sum_{i=1}^{m}\mathbb{E}[\mathrm{Reg}_{T}^{(i)}(\pi;\nu)]$. Throughout this section, the notation $\lesssim$ hides constants that may depend on $T$ and $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
The intuition for (8) and (10) is as follows: for instances that under an appropriate separation condition on $\mathcal{U}$, once the reference $b\in\mathcal{B}$ is revealed, any policy $\pi$ must incur a nontrivial coordinate-wise identification error. This lower bound on the identification error can in turn be converted into a lower bound on the associated partial regret in (10). Summing them up and optimizing over $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ then yields regret lower bounds. Approach I strictly extends the interactive Fano methods for stationary bandits outlined in Section 3.1: it allows a broader separation condition and links identification to partial regret via the post-interaction reference. It underlies our proof of Theorem 1, which establishes the logarithmic lower bound for $L\geq 3$.
However, as discussed after Theorem 1, when $L=2$ it is unclear how to apply Approach I. To overcome this, we develop a complementary method, again based on the conditional Fano inequality (4) and the post-interaction reference.
Approach II: Coordinate-Wise Information Cost. We again work under the product structure of $\mathcal{V}$, with $\mathcal{U}$ and $\mathcal{B}$ chosen as in (9) for each $i\in[m]$. Conditional on $(\mathcal{F}_{t-1},b)$, the action $a_{t}$ is drawn from the policy’s internal randomness and is therefore independent of $u$. Hence $I(u;a_{t}\mid\mathcal{F}_{t-1},b)=0$, and the chain rule gives
$I(u;Z\mid b)=\sum_{t=1}^{T}I\bigl(u;a_{t}\mid\mathcal{F}_{t-1},b\bigr)+\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr)=\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr).$
Consequently, (4) implies
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\!\bigl(\widehat{u}(Z,b)\neq u\bigr)\geq 1-\frac{\sum_{t=1}^{T}I\bigl(u;r_{t}\mid\mathcal{F}_{t-1},a_{t},b\bigr)+\log 2}{\log|\mathcal{U}|}.$ (11)
The advantage of (11) is that it yields regret lower bounds without requiring a direct reduction from small (partial) regret to correct (coordinate-wise) identification, a step needed in Approach I but one that may fail in some nonstationary constructions. Instead, we take an information-cost view: we construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ and an estimator $\widehat{u}(Z,b)$ such that for some $\rho\in(0,1)$,
[…]
An intuitive interpretation of (11) and (12) is as follows: once the reference is revealed, any policy $\pi$ must achieve a nontrivial level of coordinate-wise identification accuracy. Thus, (11) yields a lower bound on the information that must be acquired. In our construction, acquiring this information necessarily incurs partial regret. We use this approach to prove Theorem 2, which establishes the logarithmic lower bound for $L=2$ and complements Theorem 1.
Remark 1 (Comparison with the Classical KL-Based Approach).
Classical KL-based arguments also relate regret to information acquisition, but they are inherently pairwise: they compare transcript laws under two alternative environments and convert the resulting KL divergence into a regret lower bound (Bubeck et al. 2012, Garivier et al. 2019, Garivier and Moulines 2011). By contrast, our construction is genuinely multi-hypothesis, via Fano’s inequality, and is therefore better suited to capture the optimal dependence on the number of hypotheses and hence sharper regret lower bounds. Moreover, the relevant information measure in our framework is (conditional) mutual information, which is central to our analysis and can sometimes be linked to regret more directly than KL divergence. We provide a more detailed discussion of this after Theorem 2.
4 Regret Guarantees for Nonstationary Satisficing Bandits
In this section, we establish optimal satisficing regret guarantees for general nonstationary satisficing bandits. Sections 4.1 and 4.2 present the corresponding lower and upper bounds, respectively.
4.1 Lower Bounds
In the stationary setting, optimal satisficing regret is $T$-independent under realizability (Feng et al. 2024, Rouyer et al. 2024). However, as we show below, such a rate is generally impossible when change points are unknown in advance.
The Hard Instances. Consider a fixed $\Delta>0$, and let $\mathcal{E}_{L,\Delta}$ denote the class of two-armed, piecewise-stationary bandit instances with $L$ segments, where rewards are Gaussian with unit variance and the mean vector on each segment $\ell$ is either $(S+\Delta,S-\Delta)$ or $(S-\Delta,S+\Delta)$, where we recall $S$ is the satisficing threshold. In particular, the identity of the unique satisficing arm changes at every change point. Using the post-interaction reference framework of Section 3.2, we derive a sharp satisficing regret lower bound on $\mathcal{E}_{L,\Delta}$.
Theorem 1.
Suppose $L\geq 3$ and $\Delta^{2}T\geq L$. Then $\inf_{\pi\in\Pi}\sup_{\nu\in\mathcal{E}_{L,\Delta}}R_{T}^{S}(\pi;\nu)\gtrsim L\log(\Delta^{2}T/L)/\Delta.$
When $\Delta$ is a constant and $L\ll\Delta^{2}T$, Theorem 1 implies an $\Omega(L\log T)$ lower bound for general nonstationary satisficing bandits. Theorem 1 reflects the difficulty incurred by unknown change points: minimizing satisficing regret is at least as hard as localizing them, which already incurs a $T$-dependent $\Omega(L\log T)$ cost. The proof, deferred to Section 6.1, is based on Approach I introduced in Section 3.2.
As will become clear in the proof, Theorem 1 requires $L\geq 3$ because the construction (13) relies on a short stationary segment $t_{1},t_{2}]\subseteq[2,T-1]$ with $t_{2}-t_{1}\equiv c$. After the reference is revealed, one can choose two instances that differ only on two short interior intervals, whose total length $2c$ is independent of their locations. This yields the separation condition (14) and hence the reduction (10). When $L=2$, however, there is only one change point. If two instances place it at different times $\tau$ and $\tau^{\prime}$, then they disagree on the entire interval between $\tau$ and $\tau^{\prime}$, whose length grows with $|\tau-\tau^{\prime}|$ (Figure 1). It is therefore unclear how to construct a comparable separated instance family in this case and the possibility of $T$-independent satisficing regret remains.
The next theorem resolves this issue via Approach II, introduced in Section 3.2, and establishes the same scaling for all $L\geq 2$ under a slightly stronger condition relating $\Delta$, $T$, and $L$.
[…]
Figure 1: Schematic illustration: the full horizontal line represents the time horizon $T$, blue/red segments denote the two stationary environments, and the shaded region marks where the two instances disagree.
Theorem 2.
Suppose $L\geq 2$ and $\Delta^{2}T\geq 13L$. Then $\inf_{\pi\in\Pi}\sup_{\nu\in\mathcal{E}_{L,\Delta}}R_{T}^{S}(\pi;\nu)\gtrsim L\log(\Delta^{2}T/L)/\Delta.$
The full proof of Theorem 2 is deferred to Section 6.2. The theorem also offers a complementary intuition for the $\Omega(L\log T)$ scaling: regardless of the policy, one must acquire $\Omega(\log T)$ information per segment, since the reference can be used to reduce the identification error below a constant. The information-cost bound then implies that acquiring this information necessarily incurs $\Omega(\log T)$ satisficing regret per segment.
Challenge with KL-based Arguments. We also note that it is unclear whether such results can be derived using a KL-based argument. To see the difficulty, consider two instances $\nu,\nu^{\prime}$ that differ only on a disagreement region $D(\nu,\nu^{\prime})\subseteq[T]$. Under Gaussian rewards, the per-round KL divergence at time $t\in D(\nu,\nu^{\prime})$ is of order
$D_{\mathrm{KL}}\!\bigl(\mathcal{N}(\mu_{t}^{\nu}(a),1),\,\mathcal{N}(\mu_{t}^{\nu^{\prime}}(a),1)\bigr)\asymp\bigl(\mu_{t}^{\nu}(a)-\mu_{t}^{\nu^{\prime}}(a)\bigr)^{2}\asymp\Delta^{2},$
so the total transcript KL satisfies $D_{\mathrm{KL}}(P_{\nu}^{\pi},P_{\nu^{\prime}}^{\pi})\approx\Delta^{2}\,|D(\nu,\nu^{\prime})|.$ Thus, the KL divergence is driven primarily by the length of the disagreement region. In particular, it does not directly distinguish whether the samples collected there incur satisficing regret or merely provide information.
Our framework overcomes this difficulty by working instead with (conditional) mutual information, which can be upper bounded by the satisficing regret itself, so that any nontrivial identification task necessarily incurs satisficing regret.
4.2 Upper Bounds
We now present an algorithm whose satisficing regret matches the lower bounds (up to problem-dependent constants). We begin by defining the windowed statistics used throughout.
Windowed Statistics. For any buffer $\mathcal{B}$ of length $m$, define the dyadic window set $\mathrm{WinSet}(\mathcal{B}):=\{2^{k}:2^{k}\leq m,\ k\in\mathbb{N}\}$. For each $w\in\mathrm{WinSet}(\mathcal{B})$, define the average of the most recent $w$ entries in the buffer $\mathcal{B}$ by $\mathrm{AvgLast}(\mathcal{B},w):=\frac{1}{w}\sum_{s=m-w+1}^{m}\mathcal{B}[s].$ Given a confidence radius $\beta(\cdot)$, define the windowed lower and upper confidence bounds $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B}):=\max_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)-\beta(w)\big\}$ and $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})=\min_{w\in\mathrm{WinSet}(\mathcal{B})}\big\{\mathrm{AvgLast}(\mathcal{B},w)+\beta(w)\big\}.$
Input: horizon $T$, threshold $S$, arms $[K]$, $\beta(w)\ \leftarrow\ \sqrt{2\big(4\log T+\log K\big)/w}$ for any $w\geq 1.$
[…]
for $t=1,2,\dots,T$ do
 if leader=None then
      Pull $a_{t}$ as the next arm in a fixed round–robin over $[K]$, observe $r_{t}$, append to $\mathcal{B}_{a_{t}}^{(\kappa_{a_{t}})}$.
[…]
 else
      Pull $a_{t}\leftarrow$ leader, observe $r_{t}$, append to $\mathcal{B}_{a_{t}}^{(\kappa_{a_{t}})}$.
[…]
Algorithm 1 Nonstationary–Sat (algorithm for nonstationary satisficing bandits)
Algorithm 1 maintains a leader state. When $\textsf{leader}=\texttt{None}$, it explores until it certifies that an arm is satisficing. When a leader is present, the algorithm exploits this arm while continuously testing whether it remains satisficing. The key difficulty in piecewise-stationary environments is that, after a change point, stale samples can be misleading. The windowed statistics address this issue: $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})$ takes the most optimistic lower bound over dyadic suffix windows, while $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})$ takes the most pessimistic upper bound. Consequently, if an arm’s current mean exceeds $S$, then eventually $\mathrm{LCB}^{\operatorname{win}}(\mathcal{B})\geq S$; conversely, if the current leader becomes non-satisficing, then for a sufficiently recent window we have $\mathrm{UCB}^{\operatorname{win}}(\mathcal{B})<S$, triggering a reset and renewed exploration.
We now state the satisficing regret guarantee for Algorithm 1.
Theorem 3.
Let $\pi$ be Algorithm 1. Under Assumption 1, we have
$R^{S}_{T}(\pi;\nu)\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}\big(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T$
where $\Delta_{a,\ell}:=(S-\mu_{a,\ell})_{+}$ and $\Delta_{\ell}^{*}:=\mu_{\ell}^{*}-S.$
Specializing to $\nu\in\mathcal{E}_{L,\Delta}$, Theorem 3 yields an $\mathcal{O}(L\log T/\Delta)$ upper bound, matching Theorems 1 and 2 up to $T$-independent factors. Moreover, the dependence $\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}$ also appears in stationary satisficing regret bounds; see, e.g., (Michel et al. 2022, Theorem 2).
The $\log T$ factor comes from concentration. With $\beta(w)\asymp\sqrt{\log T/w}$, certifying $\mu>S$ (via $\mathrm{LCB}^{\operatorname{win}}\geq S$) or rejecting $\mu<S$ (via $\mathrm{UCB}^{\operatorname{win}}<S$) requires only $w=\mathcal{O}(\log T/\Delta^{2})$ samples, where $\Delta$ is the relevant gap to the threshold. The windowed bounds also prevent frequent resets: after a reset, if the environment remains unchanged, then with high probability a satisficing leader is not demoted and a non-satisficing candidate is not promoted. Thus, within each stationary segment, the algorithm spends only logarithmically many rounds either finding a satisficing leader or detecting that the current leader has fallen below $S$; afterward it plays a satisficing arm and incurs no further regret. The full proof of Theorem 3 is deferred to Section 6.3.
5 When Is $T$-Independent Satisficing Regret Achievable
It remains unclear which structural conditions permit $T$-independent satisficing regret in nonstationary $K$-armed bandits. A common feature of the hard instances behind the $\Omega(\log T)$ lower bounds in Section 4 is the presence of threshold crossings: an arm may be satisficing early on but later down-cross below $S$ at an unknown time, and the identity of the realizable arm may also switch across segments. Such changes force the learner to repeatedly certify satisficing arms, which incurs a logarithmic cost.
In this section, we identify a regime in which these obstacles disappear. Specifically, if (i) there exists an always-realizable arm that remains strictly satisficing throughout, and (ii) no arm ever transitions from satisficing to non-satisficing (i.e., no down-crossings), then constant satisficing regret—independent of both the horizon $T$ and the number of stationary segments LL—is achievable.
Formally, we impose the following assumptions.
Assumption 2 (Always-Realizability).
$\exists a^{*}\in[K]$ such that $\mu_{a^{*},\ell}>S$ for all $\ell\in[0,L-1]$.
Assumption 3 (No Down-Crossing).
For each arm $a\in[K]$, the indicator sequence $\bigl\{\mathds{1}\{\mu_{t}(a)\geq S\}\bigr\}_{t\in[T]}$ is nondecreasing in $t$.
Under Assumptions 2 and 3, Simple-Sat (Algorithm 1 in Michel et al. (2022)), originally designed for the stationary setting, can be applied directly in the nonstationary case and achieves constant satisficing regret. Below, $\widehat{\mu}_{a}(t)$ denotes the empirical mean reward of arm $a$ based on observations collected up to time $t$.
Input: threshold $S$, arms $[K]$.
For time steps $t=1,\dots,K$, pull arm $a_{t}=t$.
for $t=K+1,K+2,\dots$ do
[…]
Algorithm 2 Simple-Sat (Michel et al. 2022)
Theorem 4.
Let $\pi$ be Algorithm 2. Under Assumptions 2 and 3, we have
$R_{T}^{S}(\pi;\nu)\lesssim\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\Big(1+\frac{1}{\Delta_{a,\min}^{2}}+\frac{1}{\Delta_{*}^{2}}\Big),$
where $\Delta_{a,\min}:=\min_{\ell\in[0,L-1]}\{\Delta_{a,\ell}:\ \Delta_{a,\ell}>0\}$, $\Delta_{a,\max}:=\max_{\ell\in[0,L-1]}\{\Delta_{a,\ell}:\ \Delta_{a,\ell}>0\}$ and $\Delta_{*}:=\min_{\ell\in[0,L-1]}(\mu_{a^{*},\ell}-S).$
Theorem 4 establishes a constant satisficing regret bound that does not depend on $T$ or $L$, and it recovers the stationary guarantee when $L=1$; see Michel et al. (2022). The key proof idea is that, under Assumptions 2 and 3, the always-realizable arm and any non-satisficing arm can be identified in finite expected time before the latter ever becomes satisficing. The full proof is deferred to Section 6.4. We emphasize that Assumptions 2 and 3 are sufficient (but not necessary) for constant satisficing regret. Characterizing necessary and sufficient structural conditions remains an interesting direction for future work.
6 Proofs
6.1 Proof of Theorem 1
We prove Theorem 1 using Approach I in Section 3.2.
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. Assume $\xi_{t}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume that $L$ is odd and that 2T/(L1)2T/(L-1) is an integer; the even-$L$ case can be handled similarly by adding an initial change point at time $2$. Let $m:=\frac{2T}{L-1}$ and $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\}$ for any $\ell\in[(L-1)/2],$ so the horizon is partitioned into (L1)/2(L-1)/2 disjoint blocks. In each block $\ell$, there is exactly one special window of length $l$, whose start time is unknown. Fix $n,l\geq 1$ with $nl\leq m-2$, and define the candidate within-block start times by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw
$\nu=(\nu_{1},\dots,\nu_{(L-1)/2}),\qquad\nu_{\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}),$
and interpret $\nu_{\ell}$ as the start time of the special window in block $\ell$, with corresponding absolute start time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each $t\in\mathcal{I}_{\ell}$, set
$(\mu_{t}(1),\mu_{t}(2))=\begin{cases}(S-\Delta,S+\Delta),&t\in[\widetilde{\nu}_{\ell},\widetilde{\nu}_{\ell}+l-1],\\ (S+\Delta,S-\Delta),&t\in\mathcal{I}_{\ell}\setminus[\widetilde{\nu}_{\ell},\widetilde{\nu}_{\ell}+l-1].\end{cases}$ (13)
This construction yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ be the complete interaction transcript, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{(L-1)/2})$ for the collection of all components except the $\ell$-th. A direct calculation shows that for any $\ell\in[(L-1)/2]$, any fixed $\tau_{-\ell}$, and any two distinct $\tau_{\ell}\neq\tau_{\ell}^{\prime}$, we have
$D_{\mathrm{KL}}(P_{\tau_{\ell},\tau_{-\ell}}\|P_{\tau^{\prime}_{\ell},\tau_{-\ell}})=4\Delta^{2}l.$ (14)
Therefore, by (8),
$\inf_{\widehat{\nu}_{\ell}(Z,\nu_{-\ell})}\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell}\bigr)$ $\geq 1-\frac{\frac{1}{n^{(L-3)/2}n^{2}}\sum_{\tau_{-\ell}}\sum_{\tau_{\ell},\tau^{\prime}_{\ell}}D_{\mathrm{KL}}(P_{\tau_{\ell},\tau_{-\ell}}\|P_{\tau^{\prime}_{\ell},\tau_{-\ell}})+\log 2}{\log n}$
$\geq 1-\frac{4\Delta^{2}l+\log 2}{\log n}\geq\frac{1}{2},$
where the last inequality holds by choosing $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$ and taking $n$ sufficiently large.
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$ to lower bound partial regret. For each window $\tau\in\mathcal{T}$ in block $\ell$, construct the counter and the estimator:
$N_{2}^{(\ell)}(\tau):=\sum_{t=(\ell-1)m+\tau}^{(\ell-1)m+\tau+l-1}\mathds{1}\{a_{t}=2\},\qquad\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\arg\max_{\tau\in\mathcal{T}}N_{2}^{(\ell)}(\tau),$
with ties broken arbitrarily. Define the blockwise number of non-satisficing pulls: $M_{\ell}:=\sum_{t\in\mathcal{I}_{\ell}}\mathds{1}\{a_{t}\neq a_{t}^{*}\},$ where $a_{t}^{*}$ is the unique satisficing arm at time $t$. Noting that inside the true window arm $2$ is satisficing for all $l$ rounds, if $M_{\ell}\leq\frac{1}{4}l$, then $N_{2}^{(\ell)}(\nu_{\ell})\ \geq\ l-M_{\ell}\ \geq\ \frac{3}{4}l.$ Consequently, $M_{\ell}\leq\frac{1}{4}l$ implies $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})=\nu_{\ell}$, and thus
$\frac{1}{2}\leq\mathbb{P}\bigl(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell}\bigr)\leq\mathbb{P}\bigl(M_{\ell}>\tfrac{1}{4}l\bigr)$ $\leq\frac{4\mathbb{E}[M_{\ell}]}{l}\leq\frac{64\Delta^{2}\mathbb{E}[M_{\ell}]}{\log n}.$
Then, $\mathbb{E}[M_{\ell}]\geq\frac{\log n}{128\Delta^{2}}$ for every $\ell\in[(L-1)/2]$. Since each pull of a non-satisficing arm incurs satisficing regret $\Delta$, the partial regret for block $\ell$ is given by $\Delta\mathbb{E}[M_{\ell}].$ We thus have
$\mathbb{E}[R_{T}^{S}(\pi,\nu)]=\sum_{\ell\in[(L-1)/2]}\Delta\mathbb{E}[M_{\ell}]\geq\frac{(L-1)\log n}{256\Delta}.$
Optimize $n$. Recall that $nl+2\leq 2T/(L-1)$ and $l=\big\lceil\frac{\log n}{16\Delta^{2}}\big\rceil$. Ignoring the ceiling, it suffices to choose $n\geq 2$ such that $n\log n\leq 16\Delta^{2}(\frac{2T}{L-1}-2).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}\geq L$, we further have $16\Delta^{2}(\frac{2T}{L-1}-2)>4.$ The proof of Theorem 1 is therefore completed by applying Lemma 1.
Lemma 1.
Let $y\geq 4$. Then there exists an integer $x\geq 2$ such that $x\log x\leq y$ and $\log x\geq\frac{1}{2}\log y.$
By Lemma 1, we finish the proof by choosing $n$ such that $\mathbb{E}[R_{T}^{S}(\pi,\nu)]\gtrsim\frac{L\log(\Delta^{2}T/L)}{\Delta}.$
Proof.
Proof of Lemma 1 Let $x:=\lceil\sqrt{y}\rceil$. Then $x\geq\sqrt{y}$, so $\log x\geq\frac{1}{2}\log y$. It remains to show $x\log x\leq y$. Let $n:=\lfloor\sqrt{y}\rfloor$, so $n\geq 2$ and $y\in[n^{2},(n+1)^{2})$, which implies $x=\lceil\sqrt{y}\rceil\leq n+1$. Thus it suffices to prove
$(n+1)\log(n+1)\leq n^{2}\qquad\text{for all integers }n\geq 2.$ (15)
For $n=2$, (15) holds since $3\log 3\leq 4$. For $n\geq 3$, define the function $f(u):=u-1-\log(u+1)$ for $u\geq 3$. Then $f^{\prime}(u)=1-\frac{1}{u+1}=\frac{u}{u+1}>0,$ so $f$ is increasing on $3,\infty)$, and $f(n)\geq f(3)=2-\log 4>0$. Hence $\log(n+1)\leq n-1$ for all integers $n\geq 3$, and therefore $(n+1)\log(n+1)\leq(n+1)(n-1)=n^{2}-1\leq n^{2},$ proving (15). Since $y\geq n^{2}$, we conclude $x\log x\leq(n+1)\log(n+1)\leq n^{2}\leq y$. This completes the proof of Lemma 1. ∎
6.2 Proof of Theorem 2
We prove Theorem 2 using Approach II in Section 3.2.
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. Assume $\xi_{t}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, suppose $L$ is even and $2T/L\in\mathbb{N}$. Let $m:=\frac{2T}{L}$ and $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\}$ for any $\ell\in[L/2].$ Then the horizon is partitioned into $L/2$ disjoint blocks of length $m$, each containing one change point. Fix integers $n,l\geq 1$ with $nl\leq m-1$, and define $\mathcal{T}:=\{2,\,2+l,\,\dots,\,2+(n-1)l\},$ so that $\tau,\tau+l-1]$ and $\tau^{\prime},\tau^{\prime}+l-1]$ are disjoint for $\tau\neq\tau^{\prime}$. Draw
$\nu=(\nu_{1},\dots,\nu_{L/2}),\qquad\nu_{\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}),$
and interpret $\nu_{\ell}$ as the within-block change point in block $\ell$. Its absolute location is $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each $t\in\mathcal{I}_{\ell}$, set
$(\mu_{t}(1),\mu_{t}(2))=\begin{cases}(S-\Delta,S+\Delta),&t\in[(\ell-1)m+1,\widetilde{\nu}_{\ell}-1],\\ (S+\Delta,S-\Delta),&t\in[\widetilde{\nu}_{\ell},\ell m].\end{cases}$ (16)
This yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{L/2})$ for the collection of all components except the $\ell$-th. Fix $\ell\in[L/2]$ and $t\in[T]$, and condition on $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})=(f,a,x)$. Define
$p:=\mathbb{P}(\widetilde{\nu}_{\ell}\leq t\mid f,a,x)\mathds{1}\{a=2\}+\mathbb{P}(\widetilde{\nu}_{\ell}>t\mid f,a,x)\mathds{1}\{a=1\},\quad P_{+}:=\mathcal{N}(S+\Delta,1),\quad P_{-}:=\mathcal{N}(S-\Delta,1).$
Given $(f,a,x)$, for any $t\in\mathcal{I}_{\ell}$, the conditional law of $r_{t}$ under the prior on $\nu_{\ell}$ is the mixture $Q:=(1-p)P_{+}+pP_{-}$. Moreover, conditionally on $(f,a,x)$, $r_{t}$ depends on $\nu_{\ell}$ only through the indicator $\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\}$. Thus, for any $t\in\mathcal{I}_{\ell}$,
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $=I\bigl(\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x\bigr)$
$=(1-p)D_{\mathrm{KL}}\bigl(P_{+}\|Q\bigr)+pD_{\mathrm{KL}}\bigl(P_{-}\|Q\bigr).$
If $t\notin\mathcal{I}_{\ell}$, then $r_{t}$ does not depend on $\nu_{\ell}$ given $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$, and therefore
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)=0,\quad\forall t\in[T]\setminus\mathcal{I}_{\ell}.$
By convexity of $D_{\mathrm{KL}}$ in its second argument, $D_{\mathrm{KL}}(P_{+}\|Q)\leq pD_{\mathrm{KL}}(P_{+}\|P_{-})$ and $D_{\mathrm{KL}}(P_{-}\|Q)\leq(1-p)D_{\mathrm{KL}}(P_{-}\|P_{+}).$ Thus, for any $t\in\mathcal{I}_{\ell}$,
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $\leq(1-p)pD_{\mathrm{KL}}(P_{+}\|P_{-})+p(1-p)D_{\mathrm{KL}}(P_{-}\|P_{+})$
$\leq p(D_{\mathrm{KL}}(P_{+}\|P_{-})+D_{\mathrm{KL}}(P_{-}\|P_{+}))=4\Delta^{2}p.$
For any $t\in\mathcal{I}_{\ell}$, taking expectation over $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$ yields
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$ $\leq 4\Delta^{2}\mathbb{E}\Bigl[\mathds{1}\{\widetilde{\nu}_{\ell}\leq t,a_{t}=2\}+\mathds{1}\{\widetilde{\nu}_{\ell}>t,a_{t}=1\}\Bigr].$
Finally, for any $\ell\in[L/2]$,
$\sum_{t=1}^{T}I\bigl(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1},a_{t},\nu_{-\ell}\bigr)\leq 4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\widetilde{\nu}_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\widetilde{\nu}_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr].$ (17)
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$ with upper bounded identification error. Fix $\ell\in[L/2]$. For each candidate change-point $\tau\in\mathcal{T}$, define $B_{\ell,\tau}:=\bigl[(\ell-1)m+\tau,\;(\ell-1)m+\tau+l-1\bigr]$ and $N_{\ell,\tau}(Z):=\sum_{t\in B_{\ell,\tau}}\mathds{1}\{a_{t}=1\}.$ Fix an integer $c$ with $1\leq c\leq l$. For each $\tau$ such that $N_{\ell,\tau}(Z)\geq c$, let $\overline{X}_{\ell,\tau}(c)$ denote the average of the first $c$ rewards of arm $1$ observed within $B_{\ell,\tau}$. For each $\tau$ such that $N_{\ell,\tau}(Z)<c$, let $\overline{X}^{\prime}_{\ell,\tau}(l-c)$ denote the average of the first lcl-c rewards of arm $2$ observed within $B_{\ell,\tau}$. Define the estimators
$\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)\geq c\ \text{ and }\ \overline{X}_{\ell,\tau}(c)\geq S\Bigr\}$
$\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)<c\ \text{ and }\ \overline{X}_{\ell,\tau}^{\prime}(l-c)\leq S\Bigr\}$
We next state a lemma that bounds the error probabilities of the two estimators above; its proof is deferred to Section 6.2.1.
Lemma 2.
Fix integers $l,c\in\mathbb{N}$ with $1\leq c\leq l$, and $\ell\in[L/2]$. Then,
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}(N_{\ell,\nu_{\ell}}(Z)<c)+n\exp(-c\Delta^{2}/2),$
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}\big(N_{\ell,\nu_{\ell}}(Z)\geq c\big)+n\exp(-(l-c)\Delta^{2}/2).$
Construct the final estimator by flipping an independent coin $\Xi\sim\mathrm{Unif}\{0,1\}$. If $\Xi=0$, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$; otherwise, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$. Combining Lemma 2 with the choices $c=\lceil l/2\rceil$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$, we obtain
$\mathbb{P}(\widehat{\nu}_{\ell}(Z,\nu_{-\ell})\neq\nu_{\ell})\ \leq 1/2+n\exp(-(l-1)\Delta^{2}/2)\ \leq 1/2+1/n^{2}\leq 3/4,$ (18)
Combining (17) and (18), and using (11), we obtain
$\mathbb{E}[R_{T}^{S}(\pi;\nu)]=\Delta\cdot\sum_{\ell\in[L/2]}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\widetilde{\nu}_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\widetilde{\nu}_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{L}{8\Delta}\Bigl(\frac{1}{4}\log n-\log 2\Bigr).$
Optimize $n$. Recall the constraints $nl\leq 2T/L-1$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$. Ignoring the ceiling for simplicity, it suffices to choose $n\geq 2$ such that $n\log n\;\leq\;\frac{\Delta^{2}}{6}(\frac{2T}{L}-1).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}/L\geq 13$, we further have $\frac{\Delta^{2}}{6}(\frac{2T}{L}-1)\geq 13/3-1/32>4.$ By Lemma 1, we may choose $n\geq 2$ so that $\mathbb{E}\!\left[R_{T}^{S}(\pi;\nu)\right]\gtrsim\frac{L}{\Delta}\log(\Delta^{2}T/L)$. This completes the proof of Theorem 2.
6.2.1 Proof of Lemma 2
In this section, we prove the upper bound on $\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ in Lemma 2; the second bound follows by an analogous argument. Let $E:=\{N_{\ell,\nu_{\ell}}(Z)<c\}$. Then
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})\leq\mathbb{P}(E)+\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}}).$
On the event $E^{\mathrm{c}}$, the estimator $\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ can make error only if either the block at the true change point is not declared, i.e., $\overline{X}_{\ell,\nu_{\ell}}<S$; or some earlier candidate is declared first, i.e., $\overline{X}_{\ell,\tau}\geq S$ for some $\tau<\nu_{\ell}$ with $N_{\ell,\tau}(Z)\geq c$.
For the first event, since $\overline{X}_{\ell,\nu_{\ell}}\sim\mathcal{N}(S+\Delta,1/c)$ on $E^{\mathrm{c}}$, the Gaussian tail bound gives $\mathbb{P}(\overline{X}_{\ell,\nu_{\ell}}<S)=\mathbb{P}\big(\overline{X}_{\ell,\nu_{\ell}}-(S+\Delta)<-\Delta\big)\leq\exp(-c\Delta^{2}/2).$
Next, fix any $\tau<\nu_{\ell}$. On $E^{\mathrm{c}}\cap\{N_{\ell,\tau}(Z)\geq c\}$, we have $\overline{X}_{\ell,\tau}\sim\mathcal{N}(S-\Delta,1/c)$, and hence $\mathbb{P}\big(\overline{X}_{\ell,\tau}\geq S\big)=\mathbb{P}\big(\overline{X}_{\ell,\tau}-(S-\Delta)\geq\Delta\big)\leq\exp(-c\Delta^{2}/2).$
Taking a union bound over the at most n1n-1 indices $\tau\in\mathcal{T}$ with $\tau<\nu_{\ell}$, we obtain
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}})\leq(n-1)\exp(-c\Delta^{2}/2)+\exp(-c\Delta^{2}/2)=n\exp(-c\Delta^{2}/2).$
6.3 Proof of Theorem 3
By the definition of the satisficing regret in (2), we have
$R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a=1}^{K}\Delta_{a,\ell}\mathbb{P}_{\pi,\nu}(a_{t}=a)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a),$ (19)
where, in the last equality, we simplify the notation by omitting the dependence on the policy $\pi$ and the environment $\nu$. We first introduce some useful notation. For $\Delta>0$, define
$n_{\Delta}:=\min\Bigl\{n\in[T]:\sqrt{\frac{2(4\log T+\log K)}{n}}\leq\Delta/2\Bigr\}\leq c_{1}\frac{\log TK}{\Delta^{2}},$
where $c_{1}>0$ is a universal constant. Consequently, we have $\beta(n)\leq\Delta/2$ for all $n$ such that $n_{\Delta}\leq n\leq T$. Define the event $\mathcal{E}_{\operatorname{win}}$ that for every arm $a$, epoch $\kappa$, time $t\leq T$, and every dyadic window $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$ whose last $w$ samples lie entirely inside a stationary segment $\ell$, we have $\Big|\mathrm{AvgLast}\big(\mathcal{B}_{a}^{(\kappa)}(t),w\big)-\mu_{a,\ell}\Big|\ \leq\ \beta(w).$ By Hoeffding’s inequality for $1$-sub-Gaussian variables, for fixed $(a,\kappa,t,w)$ whose last $w$ samples lie in some stationary segment $\ell$,
$\mathbb{P}\left(\Big|\mathrm{AvgLast}\big(\mathcal{B}_{a}^{(\kappa)}(t),w\big)-\mu_{a,\ell}\Big|\geq\beta(w)\right)$ $\leq 2\exp\big(-\frac{w}{2}\beta(w)^{2}\big)=2\exp(-4\log T-\log K)=\frac{2}{KT^{4}}.$
Taking a union bound over all $a\in[K]$, epochs $\kappa\leq T$, times $t\leq T$, and windows $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$, and using that $|\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))|\leq T$, we obtain $\mathbb{P}(\mathcal{E}_{\operatorname{win}}^{c})\leq\frac{2}{T}.$ On the event $\mathcal{E}_{\operatorname{win}}$, the windowed confidence bounds behave as follows:
  • If $\mu_{a,\ell}=S-\Delta$ and $\mathcal{B}_{a}^{(\kappa)}$ contains a suffix window of length $w\geq 2n_{\Delta}$ entirely within segment $\ell$, then $\mathrm{UCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\leq S$. Hence any leader whose mean stays $\leq S-\Delta$ in an epoch is demoted after at most $2n_{\Delta}$ pulls within a single segment.
  • Similarly, any arm whose mean stays $\geq S+\Delta$ in an epoch is certified as satisficing once it accumulates $2n_{\Delta}$ pulls within a single segment.
  • Within a fixed segment $\ell$, any non-satisficing arm can be certified as leader at most once.
  • Within a fixed segment $\ell$, once an optimal (satisficing) arm is certified as leader, it is never demoted before the segment ends.
On the event $\mathcal{E}_{\operatorname{win}}$, the exploration length on each segment $\ell$ is at most $2Kn_{\Delta_{\ell}^{*}}$: during exploration the algorithm runs a fixed round–robin schedule, and once the optimal arm on segment $\ell$ has been pulled $2n_{\Delta_{\ell}^{*}}$ times, it is certified as satisficing and exploration terminates. Moreover, since every non-satisficing arm $a$ can be certified as leader at most once in segment $\ell$ and, once certified, is demoted after at most $2n_{\Delta_{a,\ell}}$ pulls, the regret incurred on segment $\ell$ is bounded by $\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log KT\lesssim\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T,$ up to a universal constant. Therefore, we have
$R_{T}^{S}(\pi;\nu)$ $=\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\mid\mathcal{E}_{\operatorname{win}}\Big]\mathbb{P}(\mathcal{E}_{\operatorname{win}})+\mathbb{E}_{\pi,\nu}\Big[\sum_{t=1}^{T}(S-\mu_{t}(a_{t}))_{+}\mid\mathcal{E}_{\operatorname{win}}^{c}\Big]\mathbb{P}(\mathcal{E}_{\operatorname{win}}^{c})$
$\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T+T\cdot\frac{2}{T}\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T,$
thereby completing the proof of Theorem 3.
6.4 Proof of Theorem 4
In this section, we prove Theorem 4 by adapting the proof of Theorem 1 in Michel et al. (2022) from the stationary setting to our nonstationary setting under Assumptions 2 and 3. We work with the same regret representation as in equation (19): $R_{T}^{S}(\pi;\nu)=\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a).$ Following the event decomposition used in Appendix A of Michel et al. (2022), we write, for each arm $a$ and time $t$, $\left\{a_{t}=a\right\}\subset\{t=a\}\cup\left\{a_{t}=a,Z_{t}^{c}\right\}\cup\left\{a_{t}=a,Z_{t}\right\},$ where $Z_{t}:=\left\{\forall a\in[K],\hat{\mu}_{t}(a)<S\right\}$. For the first event, we have $\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathds{1}\{t=a\}\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}.$
For each arm $a\in[K]\setminus\{a^{*}\}$, define $T(a):=\max\{t\in[T]:\mu_{t}(a)<S\}$. For $n\leq T(a)$, let $\hat{\mu}_{a,n}$ denote the empirical mean of arm $a$ based on the first $n$ pulls of $a$ that occur strictly before time $T(a)$. For the second event, Assumption 3 implies that
$\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t}^{c})$ $\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\mathbb{P}(\hat{\mu}_{a,n}(a)\geq S)$
$\leq\sum_{a:\Delta_{a,\min}>0}\Delta_{a,\max}\sum_{n=1}^{T(a)}\exp(-\frac{n\Delta_{a,\min}^{2}}{2})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{a,\min}^{2}}.$
Let $\hat{\mu}_{a^{*},n}$ denote the empirical mean of arm $a^{*}$ based on the first $n$ pulls of $a^{*}$. For the last event, a similar argument, together with Assumption 2, implies that
$\sum_{\ell=0}^{L-1}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a,Z_{t})$ $=\sum_{\ell=1}^{L}\sum_{t=T_{\ell}}^{T_{\ell+1}-1}\sum_{a:\mu_{a,\ell}<S}\Delta_{a,\ell}\mathbb{P}(a_{t}=a^{*},Z_{t})\lesssim\sum_{a:\Delta_{a,\min}>0}\frac{\Delta_{a,\max}}{\Delta_{*}^{2}}.$
Combining the three events, we complete the proof of Theorem 4.
7 Conclusion
This paper studies satisficing regret in nonstationary bandits. Our results reveal a sharp contrast with the stationary realizable setting: under nonstationarity, even a single change in the environment is enough to make $T$-independent satisficing regret unattainable in general. In particular, in the realizable piecewise-stationary setting with $L$ stationary segments, we show that the optimal regret scales as $\Theta(L\log T)$ whenever $L\geq 2$. Our lower bounds are based on a new post-interaction reference framework, together with a conditional Fano argument, that strictly extends the interactive Fano methods from stationary bandits to the nonstationary setting.
[…]